### v3

# VilBias: A Study of Bias Detection through Linguistic and Visual Cues , presenting Annotation Strategies, Evaluation, and Key Challenges

Abstract
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) opens new avenues for addressing complex challenges in multimodal content analysis, particularly in biased news detection. This study introduces ViLBias, a framework that leverages state-of-the-art LLMs and VLMs to detect linguistic and visual biases in news content. We present a multimodal dataset comprising textual content and corresponding images from diverse news sources. We propose a hybrid annotation framework that combines LLM-based annotations with human review to ensure high-quality labeling while reducing costs and enhancing scalability. Our evaluation compares the performance of state-of-the-art SLMs and LLMs for both modalities (text and images) and the results reveal that while SLMs are computationally efficient, LLMs demonstrate superior accuracy in identifying subtle framing and text-visual inconsistencies. Furthermore, empirical analysis shows that incorporating visual cues alongside textual data improves bias detection accuracy by 3–5%. This study provides a comprehensive exploration of LLMs, SLMs, and VLMs as tools for detecting multimodal biases in news content and highlights their respective strengths, limitations, and potential for future applications.
keywords:
multimodal bias detection , large language models , vision-language models ,media bias , hybrid annotation
1 Introduction
The rapid development of Large Language Models (LLMs), such as the OpenAI GPT series and Meta LLaMA, has profoundly influenced natural language processing (NLP) [1]. These models demonstrate remarkable capabilities in generating coherent text and understanding complex language patterns. More recently, researchers have begun integrating LLMs with additional modalities, such as images, audio, and video, which leads to the emergence of multimodal LLMs [2]. These systems excel at tasks requiring joint inference from textual and visual inputs, including image captioning, visual question answering, and content moderation [3].
In the realm of news media, detecting bias remains a pressing concern [4]. Media outlets frequently employ subtle linguistic framing and selective visuals to shape public perception [5]. While traditional text-based methods for bias detection identify patterns in phrasing and emphasis, they often overlook the complex interplay between textual and visual elements. In contrast, multimodal LLMs [6] can analyze multiple modalities simultaneously, offering a more holistic understanding of the data. For bias detection, these systems examine how written narratives interact with accompanying visuals to frame events, highlight specific viewpoints, or influence the dissemination of information. A working example is shown in Figure 1.
Although multimodal fake news detection has received significant attention [7, 8], bias detection remains underexplored; likely due to its subjective nature of the topic, which highlights the need for targeted research in this area. In this work, bias refers to any systematic partiality or distortion in textual or visual content that shapes how an event, individual, or group is portrayed. Such partiality manifests when one perspective is consistently favored, while other viewpoints are undermined, trivialized, or dismissed. This notion of bias is informed by multiple theoretical underpinnings in communication, psychology, and media studies, such as framing theory [9], social identity theory [10], and recent developments in computational science [11]. In multimodal contexts, these biases can be more subtle that are seen not just in written language (e.g., word choice, framing of facts) but also through images, videos, and other visual cues (e.g., cropping, staging, or emotive imagery). Consequently, identifying and categorizing such biases requires a nuanced approach that integrates both theoretical perspectives and data-driven techniques.
Figure 1: Examples of biases in news. Left: Combining image and text to amplify fear—an illustration of multimodal storytelling in conspiratorial narratives. Right: The role of visual and textual multimodality in propagating bias through a fabricated claim supported by engineered imagery. Although the text alone may not exhibit bias, the image demonstrates how multimodality can reveal potential biases and manipulative framing.
A main obstacle in multimodal bias detection is the lack of high-quality, labeled datasets for training and evaluation [12]. Expert annotations, while considered the gold standard, are both costly and time-consuming to produce. Although crowdsourcing can offer efficiency at scale, it often fails to capture the subtlety required for bias detection [13], leading to noisier datasets. Overcoming this limitation is vital for improving the accuracy and reliability of multimodal bias detection systems. The goal of this study is therefore two-fold:
  1. 1.
    To present our hybrid AI-human annotation methodology for creating a scalable, high-quality multimodal bias dataset.
  2. 2.
    To evaluate various LLMs and Vision-Language Models (VLMs) in identifying subtle biases across textual and visual content.
We introduce ViLBias (VIsual and Linguistic cues for Bias detection), a framework designed for both labeling and evaluating LLMs and VLMs on nuanced bias detection. First, we propose a hybrid annotation strategy combining an AI-based labeling process with human oversight to ensure high-quality, multimodal annotations. We then evaluate state-of-the-art (SoTA) LLMs and VLMs in performing binary bias classification on our dataset. This framework builds on our released multimodal dataset, which pairs textual content with visuals from diverse news sources. The goal is to improve label quality, address nuanced biases, and reduce annotation costs .
This study addresses the following research questions:
  1. 1.
    Does multimodality add value compared to a single modality, such as text alone?
  2. 2.
    Are Small Language Models (SLMs) [^1] The definition of SLMs varies in the literature [14]; for our purposes, SLMs are models predating the ChatGPT era that rely on an encoder-only architecture. more effective as bias classifiers, or do more recent LLMs and VLMs outperform them?
Contributions
  • In this study, we present the design and implementation of a hybrid annotation approach that integrates human expertise with LLM-based annotations. We leverage open-source LLMs as annotators and outline a structured methodology for their effective use in annotation tasks, an area that remains relatively unexplored. While some of the images and text samples are drawn from our prior work [15], this study introduces a distinct annotation framework, employing multiple LLMs to improve reliability and robustness. To support reproducibility, all related data and code are available [^2] The code is accessible at https://github.com/shainarazavi/VILBias. Data can be shared upon reasonable request.
  • We assess the performance of SoTA LLMs and VLMs in detecting biases within this multimodal dataset, comparing encoder-only SLMs with autoregressive LLMs for a bias-classification task. We also explore differences in model performance across both modalities (text and images) to identify optimal usage scenarios.
Our experiments show that incorporating visual cues alongside text improves bias detection performance by 3–5%. While SLMs perform competitively in classification tasks, LLMs demonstrate stronger generative capabilities, particularly when working with limited data. These findings underscore both the promise and the current limitations of multimodal bias detection approaches.
The remainder of this paper is organized as follows: Section 2 discusses related work in multimodal bias detection. Section 3 details our hybrid annotation methodology and the ViLBias framework. Section 4 presents experimental results, and Section LABEL:sec:conclusion concludes with insights and future directions.
2 Related Work
The study of fake news can be categorized into disinformation, misinformation [16], and, less frequently, bias news [17]. A clear distinction between biased and fake news is essential for effective media analysis. Biased news typically presents information with a particular slant, emphasizing certain viewpoints while downplaying others, whereas fake news consists of entirely fabricated content intended to mislead readers. This work primarily focuses on bias detection in news.
Research in this area often employs SLMs and LLMs, treating it as a classification task. SLMs are compact, task-specific models with fewer than 100 million parameters, whereas LLMs are larger and more versatile, typically comprising billions of parameters [1]. While there is no universal consensus on the exact distinction between SLMs and LLMs, as noted in [14], this study adopts a practical categorization: models developed before ChatGPT are referred to as SLMs, while those introduced post-ChatGPT are classified as LLMs.
Annotation Process
A significant challenge in this area of research is the scarcity of labeled data. Some studies address this issue by utilizing human annotations, often sourced from crowd workers [18], while others rely on domain experts [19] or, in rare cases, AI-generated labels [20]. The process of manual labeling typically involves human annotators with extensive knowledge, making it costly, time-consuming, and prone to errors. To mitigate these challenges, research community start exploring AI based methods that can help scale the labeling process. For example, the use of openAI GPT is seem lately for annotation of the data [20]. The goal of this work is to alleviate data scarcity in specific domains [12].
The integration of AI in data labeling processes introduces both opportunities and challenges that must be navigated with care [21]. Human oversight remains essential to ensure the quality and fairness of labeled data, as expert annotators play a crucial role in reviewing and validating AI-generated annotations. Some recent works that consider AI and human based annotations include [22, 23, 24]. Studies such as [25] and [26] relied on AI-only annotations. Hybrid SLM-LLM approaches, as seen in [27] and [28, 29], used human + AI annotations.
Table 1: LLM-based annotations in related work.
Paper Labelling Method LLM Used Problem Solved Evaluation Metric
[30] Prompts OpenAI models General, classification, text generation Human verification
[31] Prompts GPT-3.5, GPT-4 mainly, open-sourced did not work Political claim detection Accuracy, Cohen’s Kappa
[32] Prompts, explain then annotate GPT-3.5 Question answering, crowd-sourced data Accuracy, recall, MRR, Likert scale
[20] Prompts ChatGPT Classification, comparison with MTurk workers Accuracy, intercoder agreement
[21] Prompts GPT-3.5, GPT-4, LLaMA 1 Classification, comparison with MTurk workers Accuracy, intercoder agreement
[33] Prompts GPT-3.5, GPT-4 Bias detection in hate speech classification Match/mismatch with ground truth
[34] Prompts, automatic reweighting GPT, GPT-3, GPT-4 Efficiency and reliability in annotation Precision, recall, F1, std deviation
[35] Prompts GPT-4 Image captioning, text style transfer BLEU, ROUGE, BERTScore, others
[36] Prompt, weak supervision T0++, GPT-3 Text classification, spam detection Precision, recall
[37] Prompts GPT-3, InstructGPT Token classification Macro F1, accuracy
[38] Prompts GPT-3 Comparison with human annotations on NLG, NLU tasks Accuracy, ROUGE
[39] Prompts GPT-3 Comparison with human annotations on SST-2 Precision, recall, F1, time, cost
[40] Prompts GPT-3.5-turbo Translation BLEU, ROUGE
Table 2: Summary of Multimodal Models for Fake/Biased News Detection
Model Method Summary Free Dataset [^3] COCO, etc., refer to notable datasets due to space constraints.
CLIP [41] CNN, ResNet50 Predicts captions by selecting the most similar candidate based on similarity scores. Yes aYahoo, ImageNet, SUN
MiniGPT4 [42] Vicuna Describes an image based on an input image and question. Yes RefCOCO
Fuyu-8b [43] Decoder-trans Generates captions by processing images and text-based questions. Yes COCO, etc.
imageBind-LLM [44] LLaMA Handles multiple modalities to generate image descriptions. Yes COCO, etc.
MultiModal-GPT [45] CLIP and LLaMA Creates image captions and answers image-based queries. Yes COCO, etc.
Kosmos-2 [46] Method Accepts images and questions to describe objects. RefCOCO, etc.
NExT-GPT [47] ImageBlind Works with multiple modality combinations. No MosIT Data
DreamLLM [48] CLIP Capable of generating free-form content. No COCO, etc.
SPHINX [49] LLaMA 2 Provides effective image descriptions. Yes LAION-400M
BLIP-2 [50] BLIP An enhanced version of BLIP for image descriptions. Yes COCO
LLaVA [51] Vicuna Processes image and language instructions for image-based answers. Yes COCO
FakeNewsGPT4 [52] LVLM Detects news manipulation using specialized LVLM knowledge. No DGM
LEMMA [53] LVLM Enhances LVLM with external knowledge for misinformation detection. No Twitter, Fakeddit
MCNN [54] BERT, ResNet50 Combines BERT and ResNet for feature extraction to detect fake news. No D1-D4
HAMMER [55] Attn, FFN Focuses on modality interactions for misinformation detection. Yes DGM
LLMs as Annotators
Comparative analyses of multiple studies indicate that while LLMs provide cost- and time-saving benefits for classification tasks, they also encounter challenges such as lack of representativeness, bias, sensitivity to prompt variations, and a strong preference for the English language [56]. Evaluations comparing LLM annotations with those of human experts and novices reveal moderately high agreement with human experts across most quality dimensions. Additionally, incorporating LLMs as supplementary annotators can enhance inter-annotator agreement, underscoring their value in automated quality assessment tasks [13]. Frameworks that integrate LLMs into active learning loops have also been proposed to improve annotation efficiency [57]. By selecting the most informative samples for LLM annotation and employing example reweighing techniques, these frameworks enable task-specific models trained on LLM-generated labels to outperform original models, even with minimal annotated examples.
The performance of multilingual LLMs in annotating low-resource language content has been evaluated for tasks like natural language inference, sentiment analysis, and news classification. Results indicate substantial variability, with off-the-shelf multilingual LLMs that shows poor performance in certain tasks, particularly in low-resource settings [58]. The feasibility of employing LLMs as replacements for human annotators has also been investigated. Zero-shot performance across various tasks reveals that models trained with human-generated labels generally outperform those trained with LLM-generated labels. Recent work on fake news and bias detection using LLMs for annotations or evaluations is summarized in Table 1. As highlighted in the table, these annotation methods primarily focus on text-based tasks. To the best of our knowledge, this study is the first to extend such approaches to a multimodal context.
Multi-modal Bias Detection
The integration of multiple modalities in media content necessitates advanced methods for detecting bias. Research on multimodal AI emphasize the need for datasets and models that effectively address biases arising from the combination of different data types [59, 60]. SoTA approaches include CLIP-guided learning techniques for multimodal fake news detection, which leverage cross-modal capabilities to enhance accuracy by analyzing both textual and visual elements [61]. Other related works capture fine-grained connections between images and text through multimodal interaction and graph contrastive learning [62]. A detailed literature review of techniques for detecting multimodal fake and biased news, noting the relatively limited research on the latter, is summarized in Table 2.
3 Multi-Stage Annotation Methodology
3.1 Problem Definition
In this work, we focus on the task of multimodal bias detection in news media. Formally, let $\mathcal{T}$ represent the textual component (e.g., an article or caption) and $\mathcal{V}$ represent the visual component (e.g., an accompanying image). We define the problem as follows:
Input:
A pair $(\mathcal{T},\mathcal{V})$, where $\mathcal{T}$ is a textual snippet and $\mathcal{V}$ is a corresponding visual element.
Output:
A binary label $y\in\{0,1\}$ indicating the presence (1) or absence (0) of bias in the combined textual and visual content.
Here, bias is defined as a systematic partiality or distortion in how an event, individual, or group is portrayed. Such partiality may manifest when one perspective is consistently favored, while other perspectives are undermined, trivialized, or dismissed. Our focus on multimodality recognizes that biases often emerge from the interplay between textual framing (e.g., word choice, omission of details) and visual cues (e.g., cropping, emotive imagery).
Our framework, ViLBias, is depicted in Figure 2, which illustrates three main stages:
  1. 1.
    Data Sourcing and Preprocessing
  2. 2.
    Annotations via LLM/VLM and Human Validation
  3. 3.
    Evaluations
Figure 2: ViLBias Framework. The pipeline comprises three main stages: (1) Data Collection and Preprocessing, (2) Automated and Human-Annotated Labeling, and (3) Final Evaluation.
3.2 Data Compilation
We selected a range of reputable and widely recognized news organizations for data collection, guided by prior research on media bias and source credibility [5, 63]. Our chosen outlets, CNN, BBC, The New York Times, and Al Jazeera, cover a broad spectrum of geographic regions, editorial policies, and audience trust profiles. Appendix LABEL:tab:news_sources provides detailed information about these sources.
Data were collected between May 6, 2023, and September 6, 2024, capturing evolving patterns of disinformation under shifting global events. To ensure compliant and ethical data practices, we adhered to privacy regulations, complied with website terms of service, and followed established guidelines [64].
Preprocessing focused on enhancing dataset quality and relevance. We removed duplicate articles to eliminate redundancy and filtered out multimedia-based items (e.g., URLs ending in /video, /slide, or /gallery) to concentrate on text-rich content. Articles with fewer than twenty sentences were also discarded to ensure sufficient textual depth for meaningful analysis. Each retained article was paired with its first encountered image, providing a consistent multimodal reference point.
Some of the images and text samples in this study are sourced from our prior work in same project [15]. However, the labeling scheme used here differs, as it incorporates multiple LLMs for annotation, enhancing robustness and reliability.
3.3 Multiple Annotations
To efficiently annotate a large, complex dataset, we integrated automated annotators in two categories:
  1. 1.
    LLMs for textual annotations, and
  2. 2.
    VLMs for multimodal (text-image) labels.
The prompt templates used for text-based annotations are provided in Appendix B.
A key feature of our approach is a two-step voting mechanism that increases annotation reliability. First, each LLM was queried three times per item to address stochastic variation inherent in LLMs [65], and the annotation from each LLM was determined by majority vote. Next, using multiple LLMs (three, in our case), the final annotation was decided by a second majority vote across these models.
Formally, let $A_{m}$ be the set of annotations from model $m$ after three queries. The final annotation $A_{\text{final}}$ is:
$A_{\text{final}}=\arg\max_{a\in A}\sum_{m\in M}\mathbb{1}(A_{m}=a),$
where $M$ is the set of LLMs (an odd number), $A$ is the set of all possible annotations, and $A_{m}$ is determined as:
$A_{m}=\arg\max_{a\in A}\sum_{i=1}^{3}\mathbb{1}(A_{m}^{(i)}=a),$
with $A_{m}^{(i)}$ being the $i$-th query response from model $m$. The indicator function $\mathbb{1}(A_{m}^{(i)}=a)$ equals 1 if the annotation matches $a$, and 0 otherwise. This voting approach mitigates stochastic variability within and across models, ensuring robust final annotation decisions.
Human Validation
Despite automation, human oversight remains essential for validating model outputs and calibrating the annotation pipeline. We enlisted 12 reviewers, comprising three undergraduates, four master’s students, two Ph.D. holders, and three senior researchers across computer science, media studies, political science, and linguistics. Two subject-matter experts led this process, guided by bias identification protocols outlined in Appendix B. These guidelines ensured that reviewers evaluated both linguistic and visual biases as defined by the U.S. Equal Employment Opportunity Commission [66].
Reviewers assessed a stratified sample of 10,000 annotations, classifying them as Acceptable, Needs Improvement, or Incorrect, and provided written rationales where necessary. This feedback informed iterative improvements to model prompts and annotation strategies. Disagreements were resolved through consensus discussions. After refining our approach, the inter-annotator agreement (IAA), calculated using Cohen’s Kappa, between automated labels and expert judgments reached 0.72, indicating substantial alignment.
Final Dataset
The resulting dataset is organized to support bias analyses across both textual and multimodal dimensions. Each article is assigned a stable, non-identifying unique identifier derived from the first ten digits of the SHA256 hash of its headline. Core fields include:
  • outlet, headline, article text (snippet), image, URL
To capture image semantics, we use a VLM (BLIP [67]) to generate a textual image description. The final dataset contains approximately 41,412 articles, each paired with its primary image. Two key annotation fields—nlp label and nlp-image label—indicate whether the text or the combined text-image pair is likely to exhibit bias. This schema enables diverse downstream tasks, from bias detection to explanatory modeling.
Table 3 summarizes key statistics and observed patterns. Figure 3 provides sample articles and their associated images, illustrating the multimodal scope of this collection.
Table 3: Data Analysis Summary.
Metric Value
Number of Rows 40,945
Number of Columns 7
Field Names unique_id, headline, url, article_text, image, nlp_label, nlp-image_label
Top Outlets (by count) Financial Times (4895), USA TODAY (3320), CNN (3067), The Guardian (2833), Forbes (2482)
Unique Values in nlp_label 2
Distribution of nlp_label {‘Bias’: 28,281, UnBias’: 12,664}
Unique Values in nlp-image_label 2
Distribution of nlp-image_label {‘Bias’: 22,974, ‘UnBias: 17,971}
Unique Images 40,945
Figure 3: Sample headlines and their associated images. Each image corresponds to a specific news article, illustrating the multimodal structure of the dataset.
Annotation Quality
We assess the annotation quality to address the research question: How accurate are annotations generated by LLM-based systems? To evaluate this, we selected a representative sample of 1000 records from our labeled dataset. This sample size was chosen to provide a statistically significant but manageable dataset for analysis and represents a balanced distribution between instances labeled to be biased or unbiased. The selected records were independently labeled by our team, who had no prior knowledge of the existing annotations. To assess the quality of LLM-generated annotations, we calculated precision, recall, F1 score, and accuracy, metrics widely used for evaluating classification performance.
Table 4: Performance metrics based on a majority vote from the dataset and human annotated sample, evaluated using a sample of 1000 examples. Metrics include precision, recall, F1 score, and accuracy for both text and multimodal labels.
Metric Text Label (%) Multimodal Label (%)
Precision 85.0 80.0
Recall 88.0 83.0
F1 Score 86.5 81.5
Accuracy 87.0 84.0
The results presented in Table 4 showcase the strong performance of LLM-based annotations, with all metrics surpassing 80%. This underscores the effectiveness of LLMs in handling both text-based and multimodal annotation tasks. Interestingly, text labels outperformed multimodal labels, achieving higher precision (85.0% vs. 80.0%), recall (88.0% vs. 83.0%), F1 score (86.5% vs. 81.5%), and accuracy (87.0% vs. 84.0%). These results highlight that multimodal labeling poses additional challenges, likely due to the complexities of integrating textual and visual information. Nevertheless, multimodal annotations remain competitive, with an accuracy of 84.0%, demonstrating their potential to capture subtle biases and richer context across modalities. The small gap between precision and recall for both text and multimodal labels reflects balanced and consistent performance, further validating the reliability of these annotations. Overall, these findings underscore the promise of multimodal models while also emphasizing the need for continued research to improve their ability to effectively leverage interactions between text and visuals.
4 Experimental Setup
4.1 Setting
For this work, we conducted our experiments using computing setup equipped with NVIDIA A100 Tensor Core GPUs to leverage efficient parallel processing capabilities for training and inference. The software environment included Python 3.10 with essential deep learning libraries such as PyTorch 2.0, Transformers by Hugging Face for model implementation, and OpenCV for image preprocessing. The models were trained and evaluated on Ubuntu 22.04 LTS running CUDA 12.1 for optimal GPU acceleration.
4.2 LLMs as Annotators
For text-only tasks, we utilized open source models such as Llama-3-8B-Instruct, Llama-3.1-8B-Instruct [68], and Mistral-7B-Instruct-v0.3 [69]. For text-image tasks, our approach integrated LLaVa-NeXT [70], Phi-3-Vision-128k-Instruct [71], and MiniCPM-V-2_6 [72]. To handle potential conflicts, we ensured an odd number of LLMs were used as annotators, enabling effective resolution of ties. This means that when there are disagreements among the annotators, the majority vote determines the final annotation, ensuring a definitive outcome without ambiguity. Our prompts for annotation are given in Appendix B.
4.3 Baseline Models and Evaluation Metrics
To conduct a comprehensive performance evaluation, we categorized models into SLMs and LLMs. In our study, SLMs encompass transformer architectures such as BERT [73], RoBERTa [74], DistilBERT [75], GPT-2 [76], and BART [77]. We also included multimodal configurations such as XLNet paired with VGG-19, BERT with ResNet-34, and BERT integrated with CLIP.
LLMs in our framework are represented by text-only models including LLama3.2-1B (both instruct and base variants) [78], Mistral-v0.3-instruct 7B [69], and Qwen2.5-7B-Instruct [79]. Furthermore, we examined VLMs such as InstructBlipV [80], Phi-3-vision-128k-instruct [71], Pixtral-12B-2409 [81], LLaVA-1.6 [82], and Llama-3.2-11B-Vision-Instruct [78]. This selection spans from compact, task-specific designs to versatile, billion-parameter frameworks, encompassing both text-only and multimodal platforms.
For evaluation, we tested each model under zero-shot and few-shot (5-shot) scenarios , and we also fine-tuned them using our entire dataset. We utilized an 80/10/10 split for training, validation, and testing subsets, respectively, ensuring a balanced representation of classes in each. Additionally, we implemented 5-fold cross-validation to verify the models’ robustness and employed stratified sampling to maintain uniform class distribution across the folds. Our metrics for performance assessment included standard ML metrics, such as precision, recall, F1-score, accuracy, and some qualitative evaluation through human validation.
4.4 Hyperparameters
All models were implemented using PyTorch, leveraging specific configurations to optimize performance and resource utilization. Key configurations include:
  • Data Type: torch.bfloat16 was used to enable mixed-precision training, striking a balance between memory efficiency and numerical accuracy.
  • Efficient Attention Mechanism: flash attention 2 was employed (where applicable) to accelerate attention computations while reducing memory overhead, particularly for larger models.
  • Device Mapping: Models were mapped to cuda-enabled devices to utilize GPU acceleration for faster computations.
Additionally, we employed quantization techniques when fine-tuning LLMs and VLMs to reduce memory requirements and speed up inference, particularly for large-scale models. For SLMs, regular fine-tuning without quantization was applied, as these models have a smaller parameter space and are computationally less demanding.
Table 5: Key Parameters for Training, Response Generation, and Evaluation
Parameter Details
Generation Temperature: 0.2–0.4, Maximum token limit: 1024
Batch Sizes Prompt Evaluation: 100
BERT Training: 32 (15 epochs, learning rate: $2\times 10^{-5}$, weight decay: 0.01, early stopping patience: 5)
Fine-Tuning: 1 (3 epochs, learning rate: $2\times 10^{-4}$, gradient accumulation: 1, max gradient norm: 0.3)
Learning Rate Schedulers Constant with warmup ratio of 0.03
Precision Mixed precision enabled (bf16=True)
Loss Functions CrossEntropyLoss (default)
Focal Loss (alpha=1, gamma=2) for specific tasks
Evaluation Metrics Accuracy, Precision, Recall, F1 Score, Confusion Matrices (Batch Size: 32)
Quantization 4-bit precision, double quantization enabled, nf4 quantization type
Environment and Reproducibility CUDA-enabled devices, CPU fallback when necessary
Random seed set to 0
Parallel tokenization disabled (os.environ["TOKENIZERS_PARALLELISM"] = "false")
5 Results
5.1 Overall Performance Evaluation
In this experiment, we investigate the comparative classification effectiveness of SLMs versus LLMs and VLMs. Our study specifically examines model performance across different configurations and modalities. We evaluated the models in two primary settings:
  1. 1.
    Text-only classification: Using small and large language models.
  2. 2.
    Multimodal classification: Employing small and large vision-language models.
Model Config Precision Recall F1 Score Accuracy
Small Language Models (Text-Only)
BERT-base-uncased FT 87.10% 86.50% 86.80% 86.90%
DistilBERT FT 84.30% 83.40% 83.85% 85.00%
RoBERTa-base FT 89.20% 87.90% 88.54% 89.50%
GPT-2 FT 85.50% 85.00% 85.24% 85.10%
BART FT 85.80% 85.70% 85.75% 85.80%
Large Language Models (Text-Only)
LLama3.2-1B-Instruct 0-shot 72.50% 73.40% 72.95% 67.80%
5-shot 76.20% 74.00% 75.08% 71.10%
IFT 84.00% 82.30% 83.14% 82.50%
Mistral-v0.3 7B-Instruct 0-shot 68.00% 65.00% 66.47% 67.50%
5-shot 80.10% 78.90% 79.49% 77.80%
IFT 86.20% 83.40% 84.78% 80.90%
Small Vision-Language Models (Text + Image)
XLNet + VGG-19 FT 72.00% 68.20% 70.05% 77.10%
BERT + ResNet-34 FT 75.80% 71.50% 73.59% 79.40%
BERT + CLIP FT 81.30% 73.40% 77.15% 84.20%
DistilBERT + CLIP FT 68.50% 64.00% 66.18% 74.90%
RoBERTa-base + CLIP FT 84.50% 81.40% 82.90% 83.60%
Large Vision-Language Models (Text + Image)
Phi-3-Vision-128k-Instruct 0-shot 70.40% 66.00% 68.12% 69.80%
5-shot 73.20% 71.00% 72.08% 70.50%
IFT 76.80% 78.10% 77.44% 74.00%
LLaVA-v1.6-Mistral-7B 0-shot 62.50% 60.80% 61.64% 62.70%
5-shot 68.10% 67.00% 67.54% 65.20%
IFT 75.40% 74.60% 75.00% 76.10%
LLama3.2-11B-Vision-Instruct 0-shot 65.00% 68.90% 66.88% 68.40%
5-shot 73.40% 74.20% 73.80% 72.10%
IFT 80.20% 81.90% 81.04% 82.50%
Table 6: Performance of small and large language models (text-only) and vision-language models (text + image) across configurations: 0-shot, 5-shot, fine-tuning (FT), and instruction fine-tuning (IFT). Metrics are presented in percentages.
Our evaluation reveals several key insights:
Text-only models: Table 6 provides an overview of how both SLMs and LLMs perform on text classification tasks. Among the SLMs, RoBERTa-base emerged as the best-performing model after fine-tuning, achieving an F1 score of 88.54% and an accuracy of 89.50%. Other SLMs, such as BERT and DistilBERT, also achieved strong results, although their simpler architectures and fewer parameters led to slightly lower performance.
For LLMs, Llama3.2 demonstrated robust performance, closely followed by Mistral. Their results underline the importance of training methodology and data size: instruction fine-tuning (IFT) outperformed both 5-shot and 0-shot approaches, showing that larger data samples and more demonstrations improve LLM training outcomes.
Although LLMs rely on patterns in training data to predict outputs, their predictions are not inherently probabilistic in the sense of randomness [83, 65]. During tasks like text classification, both SLMs and LLMs typically select the most likely output (argmax of the predicted distribution), which makes their behavior deterministic given fixed inputs and parameters. However, in tasks like text generation, probabilistic decoding methods (e.g., sampling with temperature or top-k sampling) may be used to introduce variability. Achieving a 100% success rate remains unrealistic for both SLMs and LLMs due to factors like limited training data, inherent ambiguities in language, and the complexity of interactions among parameters. While SLMs excel in specialized tasks due to focused optimization, LLMs rely on diverse and extensive training data to fully leverage their broader capabilities. These factors contribute to the modest performance gap observed between SLMs and LLMs for certain tasks.
Overall, these findings underscore that no single model or training approach is universally optimal. Instead, performance depends on a combination of factors, including the model’s architecture, the size and diversity of the training data, and the specifics of the fine-tuning process.
Multimodal Models: Table 6 presents the comparative performance of SLMs and large VLMs for multimodal classification tasks. Among small models, the combination of BERT with CLIP achieved the highest F1 score of 77.15% and an accuracy of 84.20% under the fine-tuning (FT) configuration. This result highlights the effectiveness of combining strong text encoders with vision models for multimodal tasks.
Large models demonstrated even greater potential, particularly with IFT. LLama3.2-11B-Vision-Instruct emerged as the best-performing model in this category, achieving an F1 score of 81.04% and an accuracy of 82.50%. The improvements underscore the importance of IFT in aligning multimodal capabilities, as it consistently outperformed both 0-shot and 5-shot configurations across models.
The results also reflect performance variations across configurations—0-shot, 5-shot, fine-tuning (FT), and IFT. For instance, small vision-language models like XLNet combined with VGG-19 showed relatively modest results, with an F1 score of 70.05% and accuracy of 77.10%. In contrast, larger models such as Phi-3-Vision-128k-Instruct demonstrated substantial improvements with IFT, reaching an F1 score of 77.44% and an accuracy of 74.00%. Similarly, LLaVA-v1.6-Mistral-7B showed the greatest gain when transitioning from 0-shot to IFT, reinforcing the value of tailored fine-tuning approaches for optimizing performance.
These results highlight the trade-offs between smaller and larger models, where small models often perform better in efficiency and simplicity, while large models leverage their scale and advanced training methodologies to achieve superior performance.
5.2 Ablation Studies
To better understand which components and training configurations most significantly impact model performance, we conducted a series of ablation experiments on fine-tuned both text-only and multimodal classification tasks. Specifically, we focused on three main aspects: data size, fine-tuning techniques, and multimodal fusion strategies.
5.2.1 Ablation 1: Training Data Size
We varied the size of the training dataset for a representative LLM (LLama3.2-1B-Instruct) and a SLM (RoBERTa-base) to examine how data quantity influences performance. The results in Figure LABEL:fig:ablation1 show that LLama3.2-1B outperforms RoBERTa-base in both F1 score and accuracy across all training data fractions. Additionally, LLama3.2-1B maintains better stability as the training data is reduced, indicating its robustness in low-resource scenarios, whereas RoBERTa-base exhibits a steeper decline in performance with reduced data availability. This shows that LLMs are more stable, and instruction fine-tuning further enhances their performance and resilience to limited training data.
5.2.2 Ablation 2: Fine-Tuning Techniques (Parameter-Efficient Methods)
Fine-tuning LLMs such as LLama3.2-1B and multimodal models like LLama3.2-11B-vision-Instruct can be computationally expensive due to the need to update all parameters during training. To address this, parameter-efficient fine-tuning (PEFT) [84] methods like Low-Rank Adaptation (LoRA) and Prefix-Tuning aim to reduce computational overhead by updating only a small subset of model parameters.
Full fine-tuning (FT) updates all model parameters during training, requiring high computational and memory resources but generally achieving the best performance. LoRA introduces low-rank trainable layers to the frozen pretrained weights, allowing significant parameter savings while maintaining performance. Prefix-Tuning appends trainable prefix embeddings to the input representations, leaving the main model frozen and thus reducing computational demands.
For this experiment, we used LLama3.2-1B-Instruct for text-only tasks and LLama3.2-11B-vision-Instruct for multimodal tasks (text + image). The models were trained on 80% of the dataset, with 20% held out for evaluation. The evaluation metrics used are pecision, recall, F1 score, and accuracy. The goal was to evaluate if these parameter-efficient methods could achieve competitive performance compared to full fine-tuning.
Table 7: Performance Comparison of Fine-Tuning Techniques for LLaMA 3.2 Models.
Model Method F1 Score Accuracy Drop (%)
LLama3.2-1B (Text) Full FT 83.1% 82.5% -
LoRA 82.3% 81.8% 0.8%
Prefix-Tuning 80.7% 80.5% 2.4%
LLama3.2-11B (Text + Vision) Full FT 81.7% 83.0% -
LoRA 81.0% 82.0% 0.7%
Prefix-Tuning 80.2% 81.2% 1.5%
Table 7 shows that full fine-tuning delivers the best performance across both text-only and multimodal tasks but comes at the cost of higher memory and computational requirements. LoRA achieves competitive performance with only a 0.7-0.8% drop in F1 score and accuracy, making it a highly practical alternative for resource-constrained environments. Prefix-tuning shows slightly more degradation, with a 1.5-2.4% drop in F1 score and accuracy, but remains viable when computational resources are extremely limited. These results demonstrate that PEFT methods can achieve near-parity with full fine-tuning while significantly reducing the computational overhead, making them suitable for deployment on edge devices or in scenarios with limited training infrastructure.
5.3 Qualitative Analysis and Bias Identification
In this section, we analyze potential biases in model predictions by examining how these models reason about certain statements. We elicit the chain-of-thought from LLMs and VLMs to determine whether their reasoning (1) relies on stereotypes or over-generalizations, or (2) correctly identifies or dismisses biases where relevant. We use our fine-tuned Llama3.2 ($1$B and $11$B), Mistral-v0.3-7B, and Phi3-vision model for generating predictions.
5.3.1 Text-based Qualitative Analysis and Bias Identification
Table 8 summarizes typical examples of each model’s responses. For each input statement, the model predicts whether there is Likely or Unlikely bias, then provides a brief rationale. We focus on two main aspects:
  1. 1.
    Misclassified Bias: Cases where the model either overestimates or underestimates the presence of bias due to stereotype-based reasoning or a failure to recognize marginalized perspectives.
  2. 2.
    Correct Classification: Cases where the model appropriately recognizes or dismisses bias based on the content and context of the statement.
Model Predic Example and Reasoning Bias Observed
Llama3.2-1B Bias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “Nurses rarely contribute to surgical innovations; doctors usually lead.”
Issue: Stereotype-based reasoning. The model assumes nurses cannot be innovators, reinforcing traditional roles and overlooking individual competence.
Llama3.2-1B Unbias Input: “The committee led by women executives proposed significant policy reforms.”
Reasoning (model): “Focus on reforms; no inherent bias.”
Issue: Fails to acknowledge or explore the possibility of gender-based biases in leadership perceptions; too dismissive.
Llama3.2-11B Unbias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “There is no explicit bias. Nurses can contribute to innovative practices, and the statement acknowledges that.”
Correct Classification: No stereotype is implied; the model recognizes nurses as capable innovators.
Mistral-7B-v0.3 Bias Input: “The startup founded by women faced challenges securing venture capital funding.”
Reasoning (model): “Historical funding gaps for women-led startups indicate bias.”
Issue & Strength: Identifies systemic funding biases (correct), but may overgeneralize by implying all such startups experience the same bias.
Mistral-7B-v0.3 Unbias Input: “The rural development program focused on uplifting marginalized communities.”
Reasoning (model): “Development programs are generally well-intentioned; bias is unlikely.”
Issue: Overlooks potential systemic biases in implementation or coverage, trivializing complexities of marginalized communities.
Mistral-7B-v0.3 Unbias Input: “The community initiative led by indigenous elders improved local education outcomes.”
Reasoning (model): “Recognizes indigenous leadership positively without implying stereotypes. No bias is evident.”
Correct Classification: The model captures a constructive scenario and does not assume negative stereotypes about indigenous communities.
Table 8: Binary bias classification (Bias/Unbias) with examples of both misclassified and correctly classified statements.
Our observations in Table 8 reveal that some models (e.g., smaller Llama3.2 variants) occasionally rely on stereotypes (e.g., downplaying a nurse’s capability). Models often overlook subtle socio-cultural biases, especially when they assume that “no explicit mention of bias implies no bias”. In some cases, models correctly identify known systemic biases (e.g., women-led startups facing funding challenges) or offer neutral interpretations when appropriate.
We also observe instances where models accurately identify or dismiss bias. For example, Llama3.2-11B correctly recognizes that highlighting a nurse’s innovative contributions is not inherently biased, and Mistral-7B-v0.3 appropriately identifies systemic gender biases for women-led startups. These cases show promise that with careful fine-tuning and context-awareness prompts, models can be guided to produce balanced and accurate assessments regarding bias.
Overall, these qualitative findings help us refine the models’ training objectives, such as incorporating more nuanced real-world scenarios into the fine-tuning data and instructing models to critically assess context.
5.3.2 Multimodal Qualitative Analysis and Bias Identification
Model Input and Analysis
Llama3.2-vision-11B
Caption: “A program empowering indigenous women to advocate for their rights.”
Reasoning: The text highlights empowerment, but the image lacks visible advocacy efforts, suggesting underrepresentation.
Bias Observed: The model assumes that the absence of explicit visual markers is equal to bias, which may be an overgeneralization.
Llama3.2-vision-11B
Caption: “A collaborative community event highlighting diverse voices in science.”
Reasoning: The text emphasizes diversity, but the image lacks representation of marginalized groups, indicating bias.
Bias Observed: The model overrelies on textual implications of diversity and fails to interpret the broader visual context.
Phi-3 Vision
Caption: “A vibrant cultural festival in Africa.”
Reasoning: The image does not convey the vibrancy described in the text, suggesting bias.
Bias Observed: The model overemphasizes textual cues without fully considering the broader cultural context depicted in the image.
Phi-3 Vision
Caption: “A bustling shared street in Toronto during COVID.”
Reasoning: While the image aligns with the description, certain social cues (e.g., accessibility) are missing, implying bias.
Bias Observed: The model fails to account for nuanced social elements, such as inclusivity markers and diverse pedestrian usage.
Llama3.2-vision-11B
Caption: “Two individuals arrive at a runway for a formal event.”
Reasoning: The individuals are dressed formally and wearing sunglasses at an airport. Nothing implies a stereotype or underrepresentation.
Bias Observed: No bias observed. The model’s reasoning is correct and neutral—there is no mismatch between the caption and the image.
Phi-3 Vision
Caption: “A speaker waving to an audience at a tech conference.”
Reasoning: The person is wearing business attire and appears to be greeting viewers. The caption is consistent with the visual context.
Bias Observed: No bias observed. The model recognizes a straightforward scenario—no overgeneralization regarding the speaker’s background or role.
Table 9: Binary classification results for VLMs. Images are displayed below model names to save horizontal space while maintaining readability.
Table 9 illustrates how LLaMA3.2-vision-11B and Phi-3 Vision interpret combined textual and visual cues to classify potential bias. We identify two main tendencies:
When a caption emphasizes marginalized groups or empowerment (e.g., “A program empowering indigenous women…”), LLaMA3.2-vision-11B often infers bias if the image lacks explicit advocacy symbols (e.g., protest signs), overgeneralizing beyond what is visually provided. A similar pattern emerges when discussing “diverse voices in science”, the model interprets the absence of visibly marginalized groups as bias without considering broader context.
Phi-3 Vision sometimes focuses on literal textual descriptors (e.g., “vibrant cultural festival”) and deems any discrepancy as bias (e.g., the festival does not appear visually colorful). It may also note missing social cues (e.g., accessibility features) to imply bias, highlighting the difficulty of inferring inclusivity from a single snapshot. This can result in under-detection of subtler, systemic biases if the immediate visual–text match seems sufficient.
Despite these tendencies, we also observe correct classifications when the text clearly aligns with the image:
When the caption “A vibrant community gathering with multiple cultural performances” truly matches an image showing various cultural booths, attire, and families interacting, LLaMA3.2-vision-11B classifies bias as Unbias, noting “no sign of exclusion.” For a caption “An academic panel featuring speakers from multiple disciplines” paired with a visibly mixed group onstage, Phi-3 Vision detects “no explicit underrepresentation”, avoiding any unwarranted bias inference.
These correct examples suggest that, with clearer text–image alignment and unambiguous markers of inclusivity, both models can produce balanced, context-sensitive judgments. However, ambiguous or subtle indicators of cultural norms or social empowerment remain challenging, which indicates a need for more context-aware multimodal reasoning.
6 Discussion
6.1 Practical and Social Impact
Our findings carry both practical and social significance. By applying LLMs and VLMs, we introduce a scalable framework for detecting biases within multimodal media. This is particularly relevant for content moderation, journalism, and public policy, where identifying subtle or underrepresented viewpoints can improve accountability and foster greater transparency.
Socially, our framework empowers more equitable media practices by revealing hidden biases that shape public opinion and perpetuate stereotypes. In practical terms, this means stakeholders can systematically uncover and address harmful or skewed narratives. Integrating human–machine collaboration also allows for efficient annotations without sacrificing critical context or cultural sensitivity, lowering the chances of purely algorithmic errors.
Beyond immediate applications, this approach opens the door for adoption by educational institutions, news outlets, and regulatory authorities to raise awareness about the nuances of media bias. In the long run, its widespread use could boost trust in digital content, broaden marginalized perspectives, and support fairer decision-making processes.
6.2 Limitations
Despite promising outcomes, our work comes with certain limitations. First, although our dataset spans a diverse range of media sources, it may not fully capture global cultural or linguistic nuances, which could affect the generalizability of our results. Second, our framework prioritizes biases that are more readily detected via textual or visual cues, potentially missing more pervasive systemic or cultural biases requiring deeper contextual insight.
Additionally, dependence on SoTA language and vision-language models brings its own challenges, mainly significant computational overhead and resource requirements, which could be prohibitive for smaller organizations or those operating in lower-resource settings. While the hybrid human–machine annotation approach improves overall reliability, it is not entirely free from subjective disagreements, especially for ambiguous content [85].
Media bias is also dynamic, shifting over time. Our dataset covers the period from May 2023 to September 2024, meaning newly emerging trends or evolving patterns may fall outside its scope. Future research should aim to include a broader range of cultures, languages, and emerging media practices, as well as more granular bias categories, while exploring real-time adaptability to keep pace with rapid media developments.
7 Conclusion
In this study, we introduced ViLBias, a comprehensive framework designed to detect bias in multimodal contexts by capturing both linguistic and visual cues. Our results demonstrate that LLMs and VLMs can effectively identify subtle biases in text and images when supported by an intelligently curated dataset and a robust human–machine hybrid annotation strategy. Together, these efforts advance our goal of improving transparency in media and increasing trust in digital content. Performance metrics such as F1 score and accuracy confirm the reliability of LLM-based annotations. Furthermore, our majority-voting mechanism mitigates stochastic variations, enhancing the stability of the classification process. We also highlight the benefits of instruction-based fine-tuning and lightweight techniques like LoRA, enabling strong results even in resource-limited environments. Taken together, these contributions shows the potential of ViLBiasin bias detection in multimodal media scenarios.
Declarations
Funding: This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflict of Interest/Competing Interests: The authors declare that there are no conflicts of interest or competing interests associated with this study.
Ethics Approval and Consent to Participate: Not applicable, as the study did not involve human participants or require ethical approval.
Consent for Publication: All authors have reviewed and approved the final version of the manuscript and consent to its publication.
Data Availability: The data supporting the findings of this study are available from the corresponding author upon reasonable request.
Code Availability: The code used to implement the methods and experiments in this study is available upon publication.
Author Contributions: S.R.: Writing – original draft, Validation, Methodology, Investigation, Conceptualization, Supervision.
C.S.: Methodology, Investigation, Evaluation, Data Analysis.
E.H.: Visualization, Methodology, Evaluation, Data Analysis.
F.O.: Writing, Visualization, Methodology, Data Analysis.
M.P.: Writing – review & editing, Methodology, Evaluation.
V.C.: Methodology, Editing.
R.J.: Methodology, Evaluation.
M.L.: Data Curation.
A.Z.: Writing – review & editing, Investigation.
V.R.K.: Writing – review & editing, Investigation.
Appendix A Data Collection
The dataset includes a diverse range of news sources such as The Guardian, CBS News, ABC News, Bloomberg, The Atlantic, and Reuters. Additionally, sources like Fox News, Forbes, CNN, CNBC, and The Hill contribute to a broad spectrum of coverage. Prominent outlets such as The New York Times, Washington Post, Al Jazeera, Time, and The Daily Beast are also included. The dataset incorporates international and regional perspectives from BBC, CBC, Global News, AP News, and The Globe and Mail. Other sources range from Politico, The Economist, The Federalist, and ProPublica to Axios, Newsweek, Daily Kos, and National Review. Conservative-leaning sources such as Newsmax, OANN, Breitbart, Daily Caller, and The Washington Examiner are present, along with more liberal-leaning platforms like HuffPost, New Yorker, and NPR/PBS NewsHour. Financial and economic perspectives are represented through WSJ, FT, and CNBC, while tabloid-style and opinion-driven outlets like New York Post, Toronto Sun, and National Post provide additional viewpoints. This collection ensures a balanced mix of mainstream, independent, and alternative news coverage.
Appendix B Annotation Guidelines
Annotators were asked to follow these specific guidelines:
  1. 1.
    Accuracy: Verify whether the annotation aligns correctly with the provided input (e.g., text or text-vision pair). For text-based tasks, the annotation must correctly reflect the content of the text. For text-vision pairs, the annotation must accurately describe the relationship between the textual and visual content.
  2. 2.
    Completeness: Ensure that the annotation captures all relevant details without omitting critical information. For instance, descriptions should be neither overly brief nor excessively verbose.
  3. 3.
    Relevance: Assess whether the annotation remains within the scope of the prompt and does not include extraneous or speculative information.
  4. 4.
    Consistency: Compare annotations across similar inputs within the sample to ensure uniformity in style and language. Discrepancies in tone, terminology, or format should be flagged.
  5. 5.
    Bias Detection: Identify any systematic biases, such as stereotypes or skewed representations in text or vision-text pairs, that might emerge from the model’s outputs. Annotators were specifically instructed to consider the following bias categories:
    • Linguistic Biases: Patterns of stereotyping or unequal representation based on language usage, tone, or terminology.
    • Visual Biases: Skewed interpretations or stereotypes inferred from visual elements.
    • Gender Biases: Disparities or stereotypical roles related to gender in text or images.
    • Racial and Ethnic Biases: Representations that reflect prejudice or unequal treatment based on race or ethnicity.
    • Cultural Biases: Misinterpretations or insensitivity toward cultural norms or symbols.
    • Disability Biases: Exclusion or misrepresentation of individuals with disabilities.
    • Age Biases: Stereotyping or underrepresentation based on age.
  6. 6.
    Tie-Breaking Validation: For cases where model annotations indicate a tie in voting, verify if the chosen annotation aligns with the overall context and the majority pattern in other cases.
Annotators were instructed to record their evaluations using a structured review form, categorizing annotations as Acceptable, Needs Improvement, or Incorrect, with detailed comments explaining their reasoning. The review also included a quantitative agreement score between human judgments and model-generated annotations to measure annotation quality. The feedback collected was used to refine model prompts and guide further improvements in the annotation pipeline.
Prompt for Text-based annotation prompt = f””” Analyze the provided text for evidence of bias. Identify the presence or absence of rhetorical techniques that contribute to bias in the news content. Provide a structured response to indicate how these techniques are used to influence the narrative. Text: {text} Rhetorical Techniques to Identify: 1. Emotional Appeal: Is the language emotionally charged to evoke strong reactions favoring one side? 2. Exaggeration and Hyperbole: Are claims exaggerated or overstated to emphasize one perspective? 3. Bias and Subjectivity: Does the text reflect polarized language or framing that favors one viewpoint? 4. Repetition: Are specific words, phrases, or ideas repeated to reinforce a particular narrative? 5. Specific Word Choices: Are charged or loaded terms used to subtly or overtly indicate bias? 6. Selective Presentation: Are facts or perspectives cherry-picked to highlight one side while ignoring others? 7. Omission of Context: Are important details or broader contexts omitted to skew the narrative? 8. Appeals to Authority: Are certain experts or authorities cited selectively to bolster a specific viewpoint? 9. Framing: Is the issue presented in a way that minimizes alternatives or discredits opposing perspectives? 10. Logical Fallacies: Are flawed arguments, such as straw man, ad hominem, or false dilemmas, employed to strengthen one side? 11. Tone: Is the tone neutral, or does it reflect sarcasm, disdain, or excessive praise for one perspective? 12. Imbalance in Representation: Are opposing views underrepresented or presented in a less favorable manner? 13. Use of Statistics or Data: Are statistics manipulated or selectively presented to support one narrative? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation concise and structured, without additional explanations. ”””
Prompt for Text and image based annotation prompt = f””” Analyze the provided text and associated image for bias. Evaluate how the text and image interact to reinforce or contradict the narrative and identify techniques used to promote a specific perspective. Text: {text} Image Description: {image_description} Evaluation Criteria: 1. Emotional Appeal: Does the text or image evoke strong emotions favoring one side? 2. Exaggeration and Hyperbole: Are claims or visual elements exaggerated to dramatize one perspective? 3. Bias and Subjectivity: Does the combination of text and image indicate clear alignment with one viewpoint? 4. Repetition: Are keywords, themes, or visual motifs repeated to reinforce a particular stance? 5. Specific Word Choices or Visual Cues: Are emotionally charged words or symbolic imagery used to promote bias? 6. Selective Presentation: Are key facts highlighted while opposing information is omitted or understated? 7. Omission of Context: Is critical background information missing in the text or image captions? 8. Appeals to Authority: Are the text or visuals aligned with selectively chosen or questionable authorities? 9. Framing: Is the issue framed to promote one perspective and discredit alternatives (e.g., through cropping, captions)? 10. Logical Fallacies: Are there flawed arguments or visual manipulations (e.g., misleading captions, doctored images)? 11. Tone: Is the tone in the text or visual style neutral, critical, or excessively favorable? 12. Imbalance in Representation: Are opposing perspectives underrepresented or negatively depicted? 13. Use of Symbolism: Are symbols, colors, or motifs in the image used to manipulate perception? 14. Use of Statistics or Data: Are charts, numbers, or data selectively presented or exaggerated? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices or Visual Cues: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Symbolism: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation precise and avoid unnecessary explanations. ”””

### v5

@@
-VilBias: A Study of Bias Detection through Linguistic and Visual Cues , presenting Annotation Strategies, Evaluation, and Key Challenges
+ViLBias: Detecting and Reasoning about Bias in Multimodal Content
 Abstract
-The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) opens new avenues for addressing complex challenges in multimodal content analysis, particularly in biased news detection.
-This study introduces ViLBias, a framework that leverages state-of-the-art LLMs and VLMs to detect linguistic and visual biases in news content.
-We present a multimodal dataset comprising textual content and corresponding images from diverse news sources.
-We propose a hybrid annotation framework that combines LLM-based annotations with human review to ensure high-quality labeling while reducing costs and enhancing scalability.
-Our evaluation compares the performance of state-of-the-art SLMs and LLMs for both modalities (text and images) and the results reveal that while SLMs are computationally efficient, LLMs demonstrate superior accuracy in identifying subtle framing and text-visual inconsistencies.
-Furthermore, empirical analysis shows that incorporating visual cues alongside textual data improves bias detection accuracy by 3–5%.
-This study provides a comprehensive exploration of LLMs, SLMs, and VLMs as tools for detecting multimodal biases in news content and highlights their respective strengths, limitations, and potential for future applications.
-keywords:
-multimodal bias detection , large language models , vision-language models ,media bias , hybrid annotation
+Detecting bias in multimodal news requires models that reason over text–image pairs, not just classify text.
+In response, we present ViLBias, a VQA-style benchmark and framework for detecting and reasoning about bias in multimodal news.
+The dataset comprises 40,945 text–image pairs from diverse outlets, each annotated with a bias label and concise rationale using a two-stage LLM-as-annotator pipeline with hierarchical majority voting and human-in-the-loop validation.
+We evaluate Small Language Models (SLMs), Large Language Models (LLMs), and Vision–Language Models (VLMs) across closed-ended classification and open-ended reasoning (oVQA), and compare parameter-efficient tuning strategies.
+Results show that incorporating images alongside text improves detection accuracy by 3–5%, and that LLMs/VLMs better capture subtle framing and text–image inconsistencies than SLMs.
+Parameter-efficient methods (LoRA/QLoRA/Adapters) recover 97–99% of full fine-tuning performance with <5% trainable parameters.
+For oVQA, reasoning accuracy spans 52–79% and faithfulness 68–89%, both improved by instruction tuning; closed accuracy correlates strongly with reasoning (r = 0.91).
+ViLBias offers a scalable benchmark and strong baselines for multimodal bias detection and rationale quality.
+The data and code is available here  [^1] https://github.com/shainarazavi/VILBias .
+Keywords Multimodal, Bias Detection, LLMs, VLMs , Media Bias, Hybrid Annotation
 1 Introduction
-The rapid development of Large Language Models (LLMs), such as the OpenAI GPT series and Meta LLaMA, has profoundly influenced natural language processing (NLP) [1].
-These models demonstrate remarkable capabilities in generating coherent text and understanding complex language patterns.
-More recently, researchers have begun integrating LLMs with additional modalities, such as images, audio, and video, which leads to the emergence of multimodal LLMs [2].
-These systems excel at tasks requiring joint inference from textual and visual inputs, including image captioning, visual question answering, and content moderation [3].
-In the realm of news media, detecting bias remains a pressing concern [4].
-Media outlets frequently employ subtle linguistic framing and selective visuals to shape public perception [5].
-While traditional text-based methods for bias detection identify patterns in phrasing and emphasis, they often overlook the complex interplay between textual and visual elements.
-In contrast, multimodal LLMs [6] can analyze multiple modalities simultaneously, offering a more holistic understanding of the data.
-For bias detection, these systems examine how written narratives interact with accompanying visuals to frame events, highlight specific viewpoints, or influence the dissemination of information.
-A working example is shown in Figure 1.
-Although multimodal fake news detection has received significant attention [7, 8], bias detection remains underexplored; likely due to its subjective nature of the topic, which highlights the need for targeted research in this area.
-In this work, bias refers to any systematic partiality or distortion in textual or visual content that shapes how an event, individual, or group is portrayed.
-Such partiality manifests when one perspective is consistently favored, while other viewpoints are undermined, trivialized, or dismissed.
-This notion of bias is informed by multiple theoretical underpinnings in communication, psychology, and media studies, such as framing theory [9], social identity theory [10], and recent developments in computational science [11].
-In multimodal contexts, these biases can be more subtle that are seen not just in written language (e.g., word choice, framing of facts) but also through images, videos, and other visual cues (e.g., cropping, staging, or emotive imagery).
-Consequently, identifying and categorizing such biases requires a nuanced approach that integrates both theoretical perspectives and data-driven techniques.
-Figure 1: Examples of biases in news.
-Left: Combining image and text to amplify fear—an illustration of multimodal storytelling in conspiratorial narratives.
-Right: The role of visual and textual multimodality in propagating bias through a fabricated claim supported by engineered imagery.
-Although the text alone may not exhibit bias, the image demonstrates how multimodality can reveal potential biases and manipulative framing.
-A main obstacle in multimodal bias detection is the lack of high-quality, labeled datasets for training and evaluation [12].
-Expert annotations, while considered the gold standard, are both costly and time-consuming to produce.
-Although crowdsourcing can offer efficiency at scale, it often fails to capture the subtlety required for bias detection [13], leading to noisier datasets.
-Overcoming this limitation is vital for improving the accuracy and reliability of multimodal bias detection systems.
-The goal of this study is therefore two-fold:
+Recent advances in natural language processing (NLP) and multimodal learning, exemplified by the OpenAI GPT series [1] and the Meta LLaMA series [2], have substantially improved the analysis and interpretation of textual and visual content.
+However, despite their capacity to process large volumes of information, these systems can also inherit and amplify biases present in both modalities.
+In multimodal settings, bias extends beyond language to include visual strategies such as image selection, cropping, staging, and emotion-provoking imagery (see Figure 1).
+In this work, we define bias as directional framing in text, image, or their interaction that skews salience or interpretation (e.g., loaded wording, selective emphasis, image selection/cropping/staging, cross-modal mismatch).
+Our focus is news media, and we distinguish framing bias from misinformation.
+The topic of multimodal fake news detection has received substantial attention [3], systematic bias detection remains comparatively underexplored, partly due to its inherent subjectivity, which underscores the need for targeted research in this domain.
+Figure 1: Examples from our dataset with rationales supporting the biased label.
+Existing benchmarks for news and social media bias remain limited, with most focusing on text-only signals.
+Resources such as MBIB [4], BABE [5], and related corpora [6] emphasize textual bias.
+Multimodal datasets (e.g., NewsBag [7], MMFakeBench [8], FakeNewsNet [9]) primarily address misinformation rather than framing bias.
+Current research shows three gaps: (i) most benchmarks target fake news/misinformation rather than framing bias; when bias is considered, resources are predominantly text-only and multimodal bias benchmarks are scarce; (ii) most resources are label-only, lacking rationales or evidence, which precludes evaluation of reasoning quality; and (iii) scalable, audited annotation protocols (e.g., LLM-as-annotator with human validation) are uncommon, limiting dataset size and consistency.
+These gaps motivate benchmarks that integrate textual and visual dimensions of bias and evaluate both closed-ended classification and open-ended reasoning.
+Table 1: Bias and related multimodal ($M$) benchmarks.
+Legend: ✓=yes, ✗=no, $\circ$=partial; Mod.=modality (T=Text, I=Image); Annot.=annotation source (AI/LLM, H=Human); Rat.=rationale; MM=multimodal checks; Reason.=reasoning.
+Benchmark Mod. Focus Annot. Rat. MM chk. Reason.
+BiasLab [10] T Bias H $\circ$
+BIASsist [11] T Bias AI/LLM
+MDAM3 [12] T+I Misinfo AI+H $\circ$
+MMFakeBench [8] T+I Misinfo AI+H $\circ$
+IndiTag [13] T Bias AI/LLM $\circ$ $\circ$
+VERITE [14] T+I Misinfo H
+MBIB [4] T Bias H $\circ$
+BABE [5] T Bias H
+BASIL [15] T Bias H
+NewsBag [7] T+I Misinfo H $\circ$
+FakeNewsNet [9] T+I Misinfo AI+H $\circ$
+ViLBias (ours) T+I Bias AI+H
+We present ViLBias, a visual question answering (VQA) style, multimodal benchmark for news media bias.
+Each text–image pair is combined with a question, a label (bias presence/type), and a concise reasoning (rationale), enabling evaluation of both prediction and reasoning.
+ViLBias consists of 40k unique multimodal pairs curated from diverse real-world news sources.
+Given the scale of the data, annotations are produced through a hybrid strategy that combines LLM-based automated labeling with extensive human review to ensure quality and consistency.
+The dataset covers multiple bias categories, ranging from overt ideological framing to subtle visual manipulations such as cropping, staging, and emotive imagery.
+We evaluate a range of state-of-the-art (SoTA) small language models (SLMs)We define SLMs as encoder-only pre-ChatGPT architectures [16]., large language models (LLMs) and vision languahe models (VLMs), assessing their ability to answer multimodal bias-related questions.
+An overview of the ViLBias framework is shown in Figure 3, and its comparison with state-of-the-art (SOTA) bias-detection benchmarks is presented in Table 1.
+Contributions
+Our work can be summarized into four main contributions:
   1. 1.
-    To present our hybrid AI-human annotation methodology for creating a scalable, high-quality multimodal bias dataset.
+    We present ViLBias, a VQA-style multimodal benchmark for news media bias.
   2. 2.
-    To evaluate various LLMs and Vision-Language Models (VLMs) in identifying subtle biases across textual and visual content.
-We introduce ViLBias (VIsual and Linguistic cues for Bias detection), a framework designed for both labeling and evaluating LLMs and VLMs on nuanced bias detection.
-First, we propose a hybrid annotation strategy combining an AI-based labeling process with human oversight to ensure high-quality, multimodal annotations.
-We then evaluate state-of-the-art (SoTA) LLMs and VLMs in performing binary bias classification on our dataset.
-This framework builds on our released multimodal dataset, which pairs textual content with visuals from diverse news sources.
-The goal is to improve label quality, address nuanced biases, and reduce annotation costs .
-This study addresses the following research questions:
-  1. 1.
-    Does multimodality add value compared to a single modality, such as text alone?
-  2. 2.
-    Are Small Language Models (SLMs) [^1] The definition of SLMs varies in the literature [14]; for our purposes, SLMs are models predating the ChatGPT era that rely on an encoder-only architecture.
-more effective as bias classifiers, or do more recent LLMs and VLMs outperform them?
-Contributions
-  • In this study, we present the design and implementation of a hybrid annotation approach that integrates human expertise with LLM-based annotations.
-We leverage open-source LLMs as annotators and outline a structured methodology for their effective use in annotation tasks, an area that remains relatively unexplored.
-While some of the images and text samples are drawn from our prior work [15], this study introduces a distinct annotation framework, employing multiple LLMs to improve reliability and robustness.
-To support reproducibility, all related data and code are available [^2] The code is accessible at https://github.com/shainarazavi/VILBias.
-Data can be shared upon reasonable request.
-  • We assess the performance of SoTA LLMs and VLMs in detecting biases within this multimodal dataset, comparing encoder-only SLMs with autoregressive LLMs for a bias-classification task.
-We also explore differences in model performance across both modalities (text and images) to identify optimal usage scenarios.
-Our experiments show that incorporating visual cues alongside text improves bias detection performance by 3–5%.
-While SLMs perform competitively in classification tasks, LLMs demonstrate stronger generative capabilities, particularly when working with limited data.
-These findings underscore both the promise and the current limitations of multimodal bias detection approaches.
-The remainder of this paper is organized as follows: Section 2 discusses related work in multimodal bias detection.
-Section 3 details our hybrid annotation methodology and the ViLBias framework.
-Section 4 presents experimental results, and Section LABEL:sec:conclusion concludes with insights and future directions.
+    We introduce BiasCorpus, a dataset of 40k text–image pairs from diverse news sources, each annotated with bias labels and rationales to frame bias detection as a multimodal task.
+  3. 3.
+    We design a hybrid annotation pipeline that stabilizes outputs through multi-LLM voting and integrates expert human review to ensure high-quality annotations.
+  4. 4.
+    We conduct a systematic evaluation of SLMs, LLMs, and VLMs, highlighting their respective strengths and limitations in multimodal media bias detection.
 2 Related Work
-The study of fake news can be categorized into disinformation, misinformation [16], and, less frequently, bias news [17].
-A clear distinction between biased and fake news is essential for effective media analysis.
-Biased news typically presents information with a particular slant, emphasizing certain viewpoints while downplaying others, whereas fake news consists of entirely fabricated content intended to mislead readers.
+Figure 2: Sample headlines and their associated images. Each image corresponds to a specific news article, illustrating the multimodal structure of the dataset.
+The study of fake news can be categorized into disinformation, misinformation [17], and bias news [18].
+Biased news typically presents information with a particular slant, emphasizing certain viewpoints while downplaying others [19], whereas fake news or misinformation consists of entirely fabricated content intended to mislead readers.
 This work primarily focuses on bias detection in news.
-Research in this area often employs SLMs and LLMs, treating it as a classification task.
-SLMs are compact, task-specific models with fewer than 100 million parameters, whereas LLMs are larger and more versatile, typically comprising billions of parameters [1].
-While there is no universal consensus on the exact distinction between SLMs and LLMs, as noted in [14], this study adopts a practical categorization: models developed before ChatGPT are referred to as SLMs, while those introduced post-ChatGPT are classified as LLMs.
-Annotation Process
-A significant challenge in this area of research is the scarcity of labeled data.
-Some studies address this issue by utilizing human annotations, often sourced from crowd workers [18], while others rely on domain experts [19] or, in rare cases, AI-generated labels [20].
-The process of manual labeling typically involves human annotators with extensive knowledge, making it costly, time-consuming, and prone to errors.
-To mitigate these challenges, research community start exploring AI based methods that can help scale the labeling process.
-For example, the use of openAI GPT is seem lately for annotation of the data [20].
-The goal of this work is to alleviate data scarcity in specific domains [12].
-The integration of AI in data labeling processes introduces both opportunities and challenges that must be navigated with care [21].
-Human oversight remains essential to ensure the quality and fairness of labeled data, as expert annotators play a crucial role in reviewing and validating AI-generated annotations.
-Some recent works that consider AI and human based annotations include [22, 23, 24].
-Studies such as [25] and [26] relied on AI-only annotations.
-Hybrid SLM-LLM approaches, as seen in [27] and [28, 29], used human + AI annotations.
-Table 1: LLM-based annotations in related work.
-Paper Labelling Method LLM Used Problem Solved Evaluation Metric
-[30] Prompts OpenAI models General, classification, text generation Human verification
-[31] Prompts GPT-3.5, GPT-4 mainly, open-sourced did not work Political claim detection Accuracy, Cohen’s Kappa
-[32] Prompts, explain then annotate GPT-3.5 Question answering, crowd-sourced data Accuracy, recall, MRR, Likert scale
-[20] Prompts ChatGPT Classification, comparison with MTurk workers Accuracy, intercoder agreement
-[21] Prompts GPT-3.5, GPT-4, LLaMA 1 Classification, comparison with MTurk workers Accuracy, intercoder agreement
-[33] Prompts GPT-3.5, GPT-4 Bias detection in hate speech classification Match/mismatch with ground truth
-[34] Prompts, automatic reweighting GPT, GPT-3, GPT-4 Efficiency and reliability in annotation Precision, recall, F1, std deviation
-[35] Prompts GPT-4 Image captioning, text style transfer BLEU, ROUGE, BERTScore, others
-[36] Prompt, weak supervision T0++, GPT-3 Text classification, spam detection Precision, recall
-[37] Prompts GPT-3, InstructGPT Token classification Macro F1, accuracy
-[38] Prompts GPT-3 Comparison with human annotations on NLG, NLU tasks Accuracy, ROUGE
-[39] Prompts GPT-3 Comparison with human annotations on SST-2 Precision, recall, F1, time, cost
-[40] Prompts GPT-3.5-turbo Translation BLEU, ROUGE
-Table 2: Summary of Multimodal Models for Fake/Biased News Detection
-Model Method Summary Free Dataset [^3] COCO, etc., refer to notable datasets due to space constraints.
-CLIP [41] CNN, ResNet50 Predicts captions by selecting the most similar candidate based on similarity scores. Yes aYahoo, ImageNet, SUN
-MiniGPT4 [42] Vicuna Describes an image based on an input image and question. Yes RefCOCO
-Fuyu-8b [43] Decoder-trans Generates captions by processing images and text-based questions. Yes COCO, etc.
-imageBind-LLM [44] LLaMA Handles multiple modalities to generate image descriptions. Yes COCO, etc.
-MultiModal-GPT [45] CLIP and LLaMA Creates image captions and answers image-based queries. Yes COCO, etc.
-Kosmos-2 [46] Method Accepts images and questions to describe objects. RefCOCO, etc.
-NExT-GPT [47] ImageBlind Works with multiple modality combinations. No MosIT Data
-DreamLLM [48] CLIP Capable of generating free-form content. No COCO, etc.
-SPHINX [49] LLaMA 2 Provides effective image descriptions. Yes LAION-400M
-BLIP-2 [50] BLIP An enhanced version of BLIP for image descriptions. Yes COCO
-LLaVA [51] Vicuna Processes image and language instructions for image-based answers. Yes COCO
-FakeNewsGPT4 [52] LVLM Detects news manipulation using specialized LVLM knowledge. No DGM
-LEMMA [53] LVLM Enhances LVLM with external knowledge for misinformation detection. No Twitter, Fakeddit
-MCNN [54] BERT, ResNet50 Combines BERT and ResNet for feature extraction to detect fake news. No D1-D4
-HAMMER [55] Attn, FFN Focuses on modality interactions for misinformation detection. Yes DGM
-LLMs as Annotators
-Comparative analyses of multiple studies indicate that while LLMs provide cost- and time-saving benefits for classification tasks, they also encounter challenges such as lack of representativeness, bias, sensitivity to prompt variations, and a strong preference for the English language [56].
-Evaluations comparing LLM annotations with those of human experts and novices reveal moderately high agreement with human experts across most quality dimensions.
-Additionally, incorporating LLMs as supplementary annotators can enhance inter-annotator agreement, underscoring their value in automated quality assessment tasks [13].
-Frameworks that integrate LLMs into active learning loops have also been proposed to improve annotation efficiency [57].
-By selecting the most informative samples for LLM annotation and employing example reweighing techniques, these frameworks enable task-specific models trained on LLM-generated labels to outperform original models, even with minimal annotated examples.
-The performance of multilingual LLMs in annotating low-resource language content has been evaluated for tasks like natural language inference, sentiment analysis, and news classification.
-Results indicate substantial variability, with off-the-shelf multilingual LLMs that shows poor performance in certain tasks, particularly in low-resource settings [58].
-The feasibility of employing LLMs as replacements for human annotators has also been investigated.
-Zero-shot performance across various tasks reveals that models trained with human-generated labels generally outperform those trained with LLM-generated labels.
-Recent work on fake news and bias detection using LLMs for annotations or evaluations is summarized in Table 1.
-As highlighted in the table, these annotation methods primarily focus on text-based tasks.
-To the best of our knowledge, this study is the first to extend such approaches to a multimodal context.
-Multi-modal Bias Detection
-The integration of multiple modalities in media content necessitates advanced methods for detecting bias.
-Research on multimodal AI emphasize the need for datasets and models that effectively address biases arising from the combination of different data types [59, 60].
-SoTA approaches include CLIP-guided learning techniques for multimodal fake news detection, which leverage cross-modal capabilities to enhance accuracy by analyzing both textual and visual elements [61].
-Other related works capture fine-grained connections between images and text through multimodal interaction and graph contrastive learning [62].
-A detailed literature review of techniques for detecting multimodal fake and biased news, noting the relatively limited research on the latter, is summarized in Table 2.
-3 Multi-Stage Annotation Methodology
+Research on bias detection typically uses SLMs for classification and LLMs for open-ended reasoning.
+Prior work indicates that SLMs can be effective on overt lexical biases when fine-tuned for specific tasks [20], but they struggle with subtler forms such as framing and ideological slant, where LLMs demonstrate stronger generalization and reasoning abilities [21, 22].
+Annotation Process Multimodal bias detection suffers from limited labeled data.
+Early efforts relied on crowd workers [23] or experts [24], with some attempts at AI-generated labels [25].
+Manual labeling is costly, slow, and error-prone, prompting interest in AI-assisted approaches [26].
+Recent studies use GPT models for annotation [27], leading to hybrid frameworks that combine AI-generated labels with human verification [28], while others rely solely on AI annotations [29].
+Multimodal Bias Detection.
+Bias in news often emerges from text–image interactions, requiring models that capture cross-modal framing [30, 31].
+Approaches using CLIP-style alignment or contrastive objectives improve detection but are largely aimed at misinformation rather than bias [32, 33].
+Benchmarks reveal pitfalls such as unimodal shortcuts, where text- or image-only models perform well on multimodal tasks [14].
+Recent efforts like VLBiasBench expose demographic biases in VLMs, highlighting the need for reasoning-oriented benchmarks [34].
+Empirically, visual cues provide modest gains (approx.
+3–5%) beyond text, underscoring the value of multimodal supervision.
+In this work, we address these gaps by introducing ViLBias, a multimodal benchmark with binary labels and reasonings, annotated via an LLM-as-annotator plus human validation pipeline, and we establish baselines for both closed-ended classification and open-ended (VQA-style) models.
+Figure 3: ViLBias Framework. The pipeline comprises three main stages: (1) Data Collection and Preprocessing, (2) Automated and Human-Annotated Labeling, and (3) Final Evaluation.
+3 ViLBias Methodology
+Our proposed framework, ViLBias, is structured into three core stages, as illustrated in Figure 3.
+First, we introduce BiasCorpus, which is a systematic collection and preparation of multimodal data relevant to news bias detection.
+Second, we employ a Human-In-The-Loop (HITL) BiasAnnotation process to incorporate two step expert-driven annotations, enhancing the reliability of the dataset.
+Finally, the curated and annotated data are integrated into a comprehensive benchmarking suite designed to evaluate and analyze model performance under diverse bias scenarios.
+Together, these components provide a robust and transparent pipeline for studying bias in multiple models.
 3.1 Problem Definition
-In this work, we focus on the task of multimodal bias detection in news media.
-Formally, let $\mathcal{T}$ represent the textual component (e.g., an article or caption) and $\mathcal{V}$ represent the visual component (e.g., an accompanying image).
-We define the problem as follows:
-Input:
-A pair $(\mathcal{T},\mathcal{V})$, where $\mathcal{T}$ is a textual snippet and $\mathcal{V}$ is a corresponding visual element.
-Output:
-A binary label $y\in\{0,1\}$ indicating the presence (1) or absence (0) of bias in the combined textual and visual content.
-Here, bias is defined as a systematic partiality or distortion in how an event, individual, or group is portrayed.
-Such partiality may manifest when one perspective is consistently favored, while other perspectives are undermined, trivialized, or dismissed.
-Our focus on multimodality recognizes that biases often emerge from the interplay between textual framing (e.g., word choice, omission of details) and visual cues (e.g., cropping, emotive imagery).
-Our framework, ViLBias, is depicted in Figure 2, which illustrates three main stages:
-  1. 1.
-    Data Sourcing and Preprocessing
-  2. 2.
-    Annotations via LLM/VLM and Human Validation
-  3. 3.
-    Evaluations
-Figure 2: ViLBias Framework. The pipeline comprises three main stages: (1) Data Collection and Preprocessing, (2) Automated and Human-Annotated Labeling, and (3) Final Evaluation.
-3.2 Data Compilation
-We selected a range of reputable and widely recognized news organizations for data collection, guided by prior research on media bias and source credibility [5, 63].
-Our chosen outlets, CNN, BBC, The New York Times, and Al Jazeera, cover a broad spectrum of geographic regions, editorial policies, and audience trust profiles.
-Appendix LABEL:tab:news_sources provides detailed information about these sources.
-Data were collected between May 6, 2023, and September 6, 2024, capturing evolving patterns of disinformation under shifting global events.
-To ensure compliant and ethical data practices, we adhered to privacy regulations, complied with website terms of service, and followed established guidelines [64].
-Preprocessing focused on enhancing dataset quality and relevance.
-We removed duplicate articles to eliminate redundancy and filtered out multimedia-based items (e.g., URLs ending in /video, /slide, or /gallery) to concentrate on text-rich content.
-Articles with fewer than twenty sentences were also discarded to ensure sufficient textual depth for meaningful analysis.
-Each retained article was paired with its first encountered image, providing a consistent multimodal reference point.
-Some of the images and text samples in this study are sourced from our prior work in same project [15].
-However, the labeling scheme used here differs, as it incorporates multiple LLMs for annotation, enhancing robustness and reliability.
-3.3 Multiple Annotations
-To efficiently annotate a large, complex dataset, we integrated automated annotators in two categories:
-  1. 1.
-    LLMs for textual annotations, and
-  2. 2.
-    VLMs for multimodal (text-image) labels.
-The prompt templates used for text-based annotations are provided in Appendix B.
-A key feature of our approach is a two-step voting mechanism that increases annotation reliability.
-First, each LLM was queried three times per item to address stochastic variation inherent in LLMs [65], and the annotation from each LLM was determined by majority vote.
-Next, using multiple LLMs (three, in our case), the final annotation was decided by a second majority vote across these models.
-Formally, let $A_{m}$ be the set of annotations from model $m$ after three queries. The final annotation $A_{\text{final}}$ is:
-$A_{\text{final}}=\arg\max_{a\in A}\sum_{m\in M}\mathbb{1}(A_{m}=a),$
-where $M$ is the set of LLMs (an odd number), $A$ is the set of all possible annotations, and $A_{m}$ is determined as:
-$A_{m}=\arg\max_{a\in A}\sum_{i=1}^{3}\mathbb{1}(A_{m}^{(i)}=a),$
-with $A_{m}^{(i)}$ being the $i$-th query response from model $m$.
-The indicator function $\mathbb{1}(A_{m}^{(i)}=a)$ equals 1 if the annotation matches $a$, and 0 otherwise.
-This voting approach mitigates stochastic variability within and across models, ensuring robust final annotation decisions.
-Human Validation
-Despite automation, human oversight remains essential for validating model outputs and calibrating the annotation pipeline.
-We enlisted 12 reviewers, comprising three undergraduates, four master’s students, two Ph.D.
-holders, and three senior researchers across computer science, media studies, political science, and linguistics.
-Two subject-matter experts led this process, guided by bias identification protocols outlined in Appendix B.
-These guidelines ensured that reviewers evaluated both linguistic and visual biases as defined by the U.S.
+The main goal of this work is to assess the ability of language models, specifically VLMs, to detect and reason about media bias.
+We introduce BiasCorpus, a dataset usable in both classification (close-ended) and VQA-style (open-ended) settings.
+In the latter, each text–image pair is paired with a query such as “Does this content exhibit bias?
+Please explain your reasoning”.
+Formally, given a textual snippet $\mathcal{T}$, a corresponding visual element $\mathcal{V}$, and a bias-related query $\mathcal{Q}$, the task is defined as
+$f:\;(\mathcal{T},\mathcal{V},\mathcal{Q})\;\mapsto\;(\mathcal{A}_{\mathrm{cls}},\mathcal{A}_{\mathrm{reasoning}}),$
+where $\mathcal{A}_{\mathrm{cls}}\in\{\text{biased},\text{not biased}\}$ is the categorical prediction and $\mathcal{A}_{\mathrm{reasoning}}$ is an open-ended explanation that justifies the decision by pointing to textual framing (e.g., emotive wording, selective omission) and/or visual cues (e.g., cropping, staging, image sentiment).
+This joint formulation allows evaluation of both predictive accuracy and reasoning quality.
+3.2 BiasCorpus Curation
+We constructed BiasCorpus by ingesting articles from Google News RSS and official outlet feeds (May 2023–September 2024), de-duplicating by canonical URLs and filtering non-article formats.
+The sources, spanning CNN, BBC, The New York Times, Al Jazeera, and others as seen in Table 2, were selected to cover varied geographies, editorial orientations, and audience trust profiles [35, 36].
+To ensure quality, we removed duplicates, discarded short articles (fewer than twenty sentences), and filtered out multimedia-only items (e.g., videos, slideshows).
+Each retained article was paired with its first encountered image to establish a consistent multimodal reference.
+Portions of the text–image pairs overlap with our prior work Anonymous due to double blind policy., but the present study introduces a distinct labeling protocol based on multi-LLM annotation with human review for robustness.
+Sample headlines and images are shown in Figure 3.
+3.3 HITL BiasAnnotation
+To efficiently annotate a large, complex dataset, we integrated automated annotators, using LLM-as-annotators for bulk labeling of text–image pairs.
+The use of LLMs for annotation has recently been shown to be an effective and scalable strategy [37].
+The prompt template is provided in Appendix C.
+A distinct aspect of our annotation pipeline is a two-step voting mechanism designed to maximize reliability while reducing stochastic variability in LLM outputs.
+The process, as shown in Figure 4 operates as follows:
+(1) Step 1 – Stabilizing each LLM: For each model, the same item is queried three times.
+Since LLMs are inherently stochastic [38], the responses may differ.
+A majority vote is applied over the three responses to obtain one stable annotation per model.
+(2) Step 2 – Consensus across models: After Step 1, each LLM contributes a single annotation.
+With three LLMs in total, a second majority vote is then applied across their outputs, yielding the final annotation for that item.
+In short, each LLM first “settles on its own answer” by voting across three attempts, and then the models collectively “agree” through another vote.
+This hierarchical voting scheme ensures robustness both within and across models, providing reliable annotations.
+Figure 4: LLM-As-Annotators with Human-In-The-Loop (HITL) Bias Annotation Framework.
+Each LLM is queried three times, and its responses are majority-voted to remove stochasticity.
+These majority-voted labels are taken from three LLMs and then majority-voted (across LLMs) again to produce a final label, which is reviewed—along with LLM reasoning—by human annotators.
+Human-In-The-Loop
+Human oversight remained essential for validating automated outputs and refining the annotation pipeline.
+We engaged 12 reviewers (three undergraduates, four master’s students, two Ph.D.
+holders, and three senior researchers) spanning computer science, media studies, political science, and linguistics.
+Two subject-matter experts led the process using bias identification protocols (Appendix B), ensuring coverage of both linguistic and visual biases as defined by the U.S.
 Equal Employment Opportunity Commission [39].
-Reviewers assessed a stratified sample of 10,000 annotations, classifying them as Acceptable, Needs Improvement, or Incorrect, and provided written rationales where necessary.
-This feedback informed iterative improvements to model prompts and annotation strategies.
-Disagreements were resolved through consensus discussions.
-After refining our approach, the inter-annotator agreement (IAA), calculated using Cohen’s Kappa, between automated labels and expert judgments reached 0.72, indicating substantial alignment.
-Final Dataset
-The resulting dataset is organized to support bias analyses across both textual and multimodal dimensions.
-Each article is assigned a stable, non-identifying unique identifier derived from the first ten digits of the SHA256 hash of its headline.
-Core fields include:
-  • outlet, headline, article text (snippet), image, URL
-To capture image semantics, we use a VLM (BLIP [67]) to generate a textual image description.
-The final dataset contains approximately 41,412 articles, each paired with its primary image.
-Two key annotation fields—nlp label and nlp-image label—indicate whether the text or the combined text-image pair is likely to exhibit bias.
-This schema enables diverse downstream tasks, from bias detection to explanatory modeling.
-Table 3 summarizes key statistics and observed patterns. Figure 3 provides sample articles and their associated images, illustrating the multimodal scope of this collection.
-Table 3: Data Analysis Summary.
+Reviewers assessed all annotations, classifying them as Acceptable, Needs Improvement, or Incorrect, and provided rationales where needed.
+Their feedback informed iterative prompt and pipeline adjustments, with disagreements resolved through consensus.
+After refinement, the inter-annotator agreement (IAA) between automated labels and expert judgments, measured with Cohen’s Kappa, reached 0.72, indicating substantial alignment.
+Table 2 summarizes dataset schema, key statistics, and observed patterns.
+3.4 Benchmarking Suite
+We evaluate models on both closed-ended classification and open-ended reasoning tasks across three model families (LLMs, SLMs and VLMs) and two accessibility settings (open-ended and closed-ended).
+First, we consider open-source SLMs as text-only baselines, including encoder-only BERT-style transformers, CLIP-style joint encoders, decoder-only models such as GPT-2, and encoder–decoder models such as BART-base, all fine-tuned on our dataset.
+Second, we examine autoregressive decoder-only LLMs, tested under zero-shot, few-shot, and instruction fine-tuning (IFT) regimes.
+In addition to open-source models, we also evaluate closed-source APIs such as GPT-4o [40] and Gemini 1.5 Pro [41].
+Third, we extend evaluation to VLMs, which are assessed not only on classification accuracy but also on their ability to generate faithful, open-ended rationales grounded in both textual and visual evidence.
+For classification tasks, we report standard metrics including precision, recall, F1 score, and accuracy.
+For open-ended reasoning, we employ an LLM-as-judge [42] framework to evaluate the quality, faithfulness, and consistency of generated rationales.
+Prompts for annotation are in Appendix C and for LLM-as-judge evaluation in Figure A.1.
+Table 2: BiasCorpus analysis summary.
 Metric Value
-Number of Rows 40,945
-Number of Columns 7
-Field Names unique_id, headline, url, article_text, image, nlp_label, nlp-image_label
-Top Outlets (by count) Financial Times (4895), USA TODAY (3320), CNN (3067), The Guardian (2833), Forbes (2482)
-Unique Values in nlp_label 2
-Distribution of nlp_label {‘Bias’: 28,281, UnBias’: 12,664}
-Unique Values in nlp-image_label 2
-Distribution of nlp-image_label {‘Bias’: 22,974, ‘UnBias: 17,971}
+Rows 40,945
+Field names unique_id, headline, url, article_text, image, multimodal_label, reasoning
+Top outlets Financial Times (4,895), USA TODAY (3,320), CNN (3,067), The Guardian (2,833), Forbes (2,482)
+Label classes 2 (Biased=22,974; Not Biased=17,971)
 Unique images 40,945
-Figure 3: Sample headlines and their associated images. Each image corresponds to a specific news article, illustrating the multimodal structure of the dataset.
-Annotation Quality
-We assess the annotation quality to address the research question: How accurate are annotations generated by LLM-based systems?
-To evaluate this, we selected a representative sample of 1000 records from our labeled dataset.
-This sample size was chosen to provide a statistically significant but manageable dataset for analysis and represents a balanced distribution between instances labeled to be biased or unbiased.
-The selected records were independently labeled by our team, who had no prior knowledge of the existing annotations.
-To assess the quality of LLM-generated annotations, we calculated precision, recall, F1 score, and accuracy, metrics widely used for evaluating classification performance.
-Table 4: Performance metrics based on a majority vote from the dataset and human annotated sample, evaluated using a sample of 1000 examples.
-Metrics include precision, recall, F1 score, and accuracy for both text and multimodal labels.
-Metric Text Label (%) Multimodal Label (%)
-Precision 85.0 80.0
-Recall 88.0 83.0
-F1 Score 86.5 81.5
-Accuracy 87.0 84.0
-The results presented in Table 4 showcase the strong performance of LLM-based annotations, with all metrics surpassing 80%.
-This underscores the effectiveness of LLMs in handling both text-based and multimodal annotation tasks.
-Interestingly, text labels outperformed multimodal labels, achieving higher precision (85.0% vs.
-80.0%), recall (88.0% vs.
-83.0%), F1 score (86.5% vs.
-81.5%), and accuracy (87.0% vs.
-84.0%).
-These results highlight that multimodal labeling poses additional challenges, likely due to the complexities of integrating textual and visual information.
-Nevertheless, multimodal annotations remain competitive, with an accuracy of 84.0%, demonstrating their potential to capture subtle biases and richer context across modalities.
-The small gap between precision and recall for both text and multimodal labels reflects balanced and consistent performance, further validating the reliability of these annotations.
-Overall, these findings underscore the promise of multimodal models while also emphasizing the need for continued research to improve their ability to effectively leverage interactions between text and visuals.
-4 Experimental Setup
-4.1 Setting
-For this work, we conducted our experiments using computing setup equipped with NVIDIA A100 Tensor Core GPUs to leverage efficient parallel processing capabilities for training and inference.
-The software environment included Python 3.10 with essential deep learning libraries such as PyTorch 2.0, Transformers by Hugging Face for model implementation, and OpenCV for image preprocessing.
-The models were trained and evaluated on Ubuntu 22.04 LTS running CUDA 12.1 for optimal GPU acceleration.
-4.2 LLMs as Annotators
-For text-only tasks, we utilized open source models such as Llama-3-8B-Instruct, Llama-3.1-8B-Instruct [68], and Mistral-7B-Instruct-v0.3 [69].
-For text-image tasks, our approach integrated LLaVa-NeXT [70], Phi-3-Vision-128k-Instruct [71], and MiniCPM-V-2_6 [72].
-To handle potential conflicts, we ensured an odd number of LLMs were used as annotators, enabling effective resolution of ties.
-This means that when there are disagreements among the annotators, the majority vote determines the final annotation, ensuring a definitive outcome without ambiguity.
-Our prompts for annotation are given in Appendix B.
-4.3 Baseline Models and Evaluation Metrics
-To conduct a comprehensive performance evaluation, we categorized models into SLMs and LLMs.
-In our study, SLMs encompass transformer architectures such as BERT [73], RoBERTa [74], DistilBERT [75], GPT-2 [76], and BART [77].
-We also included multimodal configurations such as XLNet paired with VGG-19, BERT with ResNet-34, and BERT integrated with CLIP.
-LLMs in our framework are represented by text-only models including LLama3.2-1B (both instruct and base variants) [78], Mistral-v0.3-instruct 7B [69], and Qwen2.5-7B-Instruct [79].
-Furthermore, we examined VLMs such as InstructBlipV [80], Phi-3-vision-128k-instruct [71], Pixtral-12B-2409 [81], LLaVA-1.6 [82], and Llama-3.2-11B-Vision-Instruct [78].
-This selection spans from compact, task-specific designs to versatile, billion-parameter frameworks, encompassing both text-only and multimodal platforms.
-For evaluation, we tested each model under zero-shot and few-shot (5-shot) scenarios , and we also fine-tuned them using our entire dataset.
-We utilized an 80/10/10 split for training, validation, and testing subsets, respectively, ensuring a balanced representation of classes in each.
-Additionally, we implemented 5-fold cross-validation to verify the models’ robustness and employed stratified sampling to maintain uniform class distribution across the folds.
-Our metrics for performance assessment included standard ML metrics, such as precision, recall, F1-score, accuracy, and some qualitative evaluation through human validation.
-4.4 Hyperparameters
-All models were implemented using PyTorch, leveraging specific configurations to optimize performance and resource utilization. Key configurations include:
-  • Data Type: torch.bfloat16 was used to enable mixed-precision training, striking a balance between memory efficiency and numerical accuracy.
-  • Efficient Attention Mechanism: flash attention 2 was employed (where applicable) to accelerate attention computations while reducing memory overhead, particularly for larger models.
-  • Device Mapping: Models were mapped to cuda-enabled devices to utilize GPU acceleration for faster computations.
-Additionally, we employed quantization techniques when fine-tuning LLMs and VLMs to reduce memory requirements and speed up inference, particularly for large-scale models.
-For SLMs, regular fine-tuning without quantization was applied, as these models have a smaller parameter space and are computationally less demanding.
-Table 5: Key Parameters for Training, Response Generation, and Evaluation
-Parameter Details
-Generation Temperature: 0.2–0.4, Maximum token limit: 1024
-Batch Sizes Prompt Evaluation: 100
-BERT Training: 32 (15 epochs, learning rate: $2\times 10^{-5}$, weight decay: 0.01, early stopping patience: 5)
-Fine-Tuning: 1 (3 epochs, learning rate: $2\times 10^{-4}$, gradient accumulation: 1, max gradient norm: 0.3)
-Learning Rate Schedulers Constant with warmup ratio of 0.03
-Precision Mixed precision enabled (bf16=True)
-Loss Functions CrossEntropyLoss (default)
-Focal Loss (alpha=1, gamma=2) for specific tasks
-Evaluation Metrics Accuracy, Precision, Recall, F1 Score, Confusion Matrices (Batch Size: 32)
-Quantization 4-bit precision, double quantization enabled, nf4 quantization type
-Environment and Reproducibility CUDA-enabled devices, CPU fallback when necessary
-Random seed set to 0
-Parallel tokenization disabled (os.environ["TOKENIZERS_PARALLELISM"] = "false")
-5 Results
-5.1 Overall Performance Evaluation
-In this experiment, we investigate the comparative classification effectiveness of SLMs versus LLMs and VLMs.
-Our study specifically examines model performance across different configurations and modalities.
-We evaluated the models in two primary settings:
-  1. 1.
-    Text-only classification: Using small and large language models.
-  2. 2.
-    Multimodal classification: Employing small and large vision-language models.
-Model Config Precision Recall F1 Score Accuracy
-Small Language Models (Text-Only)
-BERT-base-uncased FT 87.10% 86.50% 86.80% 86.90%
-DistilBERT FT 84.30% 83.40% 83.85% 85.00%
-RoBERTa-base FT 89.20% 87.90% 88.54% 89.50%
-GPT-2 FT 85.50% 85.00% 85.24% 85.10%
-BART FT 85.80% 85.70% 85.75% 85.80%
-Large Language Models (Text-Only)
-LLama3.2-1B-Instruct 0-shot 72.50% 73.40% 72.95% 67.80%
-5-shot 76.20% 74.00% 75.08% 71.10%
-IFT 84.00% 82.30% 83.14% 82.50%
-Mistral-v0.3 7B-Instruct 0-shot 68.00% 65.00% 66.47% 67.50%
-5-shot 80.10% 78.90% 79.49% 77.80%
-IFT 86.20% 83.40% 84.78% 80.90%
-Small Vision-Language Models (Text + Image)
-XLNet + VGG-19 FT 72.00% 68.20% 70.05% 77.10%
-BERT + ResNet-34 FT 75.80% 71.50% 73.59% 79.40%
-BERT + CLIP FT 81.30% 73.40% 77.15% 84.20%
-DistilBERT + CLIP FT 68.50% 64.00% 66.18% 74.90%
-RoBERTa-base + CLIP FT 84.50% 81.40% 82.90% 83.60%
-Large Vision-Language Models (Text + Image)
-Phi-3-Vision-128k-Instruct 0-shot 70.40% 66.00% 68.12% 69.80%
-5-shot 73.20% 71.00% 72.08% 70.50%
-IFT 76.80% 78.10% 77.44% 74.00%
-LLaVA-v1.6-Mistral-7B 0-shot 62.50% 60.80% 61.64% 62.70%
-5-shot 68.10% 67.00% 67.54% 65.20%
-IFT 75.40% 74.60% 75.00% 76.10%
-LLama3.2-11B-Vision-Instruct 0-shot 65.00% 68.90% 66.88% 68.40%
-5-shot 73.40% 74.20% 73.80% 72.10%
-IFT 80.20% 81.90% 81.04% 82.50%
-Table 6: Performance of small and large language models (text-only) and vision-language models (text + image) across configurations: 0-shot, 5-shot, fine-tuning (FT), and instruction fine-tuning (IFT).
-Metrics are presented in percentages.
-Our evaluation reveals several key insights:
-Text-only models: Table 6 provides an overview of how both SLMs and LLMs perform on text classification tasks.
-Among the SLMs, RoBERTa-base emerged as the best-performing model after fine-tuning, achieving an F1 score of 88.54% and an accuracy of 89.50%.
-Other SLMs, such as BERT and DistilBERT, also achieved strong results, although their simpler architectures and fewer parameters led to slightly lower performance.
-For LLMs, Llama3.2 demonstrated robust performance, closely followed by Mistral.
-Their results underline the importance of training methodology and data size: instruction fine-tuning (IFT) outperformed both 5-shot and 0-shot approaches, showing that larger data samples and more demonstrations improve LLM training outcomes.
-Although LLMs rely on patterns in training data to predict outputs, their predictions are not inherently probabilistic in the sense of randomness [83, 65].
-During tasks like text classification, both SLMs and LLMs typically select the most likely output (argmax of the predicted distribution), which makes their behavior deterministic given fixed inputs and parameters.
-However, in tasks like text generation, probabilistic decoding methods (e.g., sampling with temperature or top-k sampling) may be used to introduce variability.
-Achieving a 100% success rate remains unrealistic for both SLMs and LLMs due to factors like limited training data, inherent ambiguities in language, and the complexity of interactions among parameters.
-While SLMs excel in specialized tasks due to focused optimization, LLMs rely on diverse and extensive training data to fully leverage their broader capabilities.
-These factors contribute to the modest performance gap observed between SLMs and LLMs for certain tasks.
-Overall, these findings underscore that no single model or training approach is universally optimal.
-Instead, performance depends on a combination of factors, including the model’s architecture, the size and diversity of the training data, and the specifics of the fine-tuning process.
-Multimodal Models: Table 6 presents the comparative performance of SLMs and large VLMs for multimodal classification tasks.
-Among small models, the combination of BERT with CLIP achieved the highest F1 score of 77.15% and an accuracy of 84.20% under the fine-tuning (FT) configuration.
-This result highlights the effectiveness of combining strong text encoders with vision models for multimodal tasks.
-Large models demonstrated even greater potential, particularly with IFT.
-LLama3.2-11B-Vision-Instruct emerged as the best-performing model in this category, achieving an F1 score of 81.04% and an accuracy of 82.50%.
-The improvements underscore the importance of IFT in aligning multimodal capabilities, as it consistently outperformed both 0-shot and 5-shot configurations across models.
-The results also reflect performance variations across configurations—0-shot, 5-shot, fine-tuning (FT), and IFT.
-For instance, small vision-language models like XLNet combined with VGG-19 showed relatively modest results, with an F1 score of 70.05% and accuracy of 77.10%.
-In contrast, larger models such as Phi-3-Vision-128k-Instruct demonstrated substantial improvements with IFT, reaching an F1 score of 77.44% and an accuracy of 74.00%.
-Similarly, LLaVA-v1.6-Mistral-7B showed the greatest gain when transitioning from 0-shot to IFT, reinforcing the value of tailored fine-tuning approaches for optimizing performance.
-These results highlight the trade-offs between smaller and larger models, where small models often perform better in efficiency and simplicity, while large models leverage their scale and advanced training methodologies to achieve superior performance.
-5.2 Ablation Studies
-To better understand which components and training configurations most significantly impact model performance, we conducted a series of ablation experiments on fine-tuned both text-only and multimodal classification tasks.
-Specifically, we focused on three main aspects: data size, fine-tuning techniques, and multimodal fusion strategies.
-5.2.1 Ablation 1: Training Data Size
-We varied the size of the training dataset for a representative LLM (LLama3.2-1B-Instruct) and a SLM (RoBERTa-base) to examine how data quantity influences performance.
-The results in Figure LABEL:fig:ablation1 show that LLama3.2-1B outperforms RoBERTa-base in both F1 score and accuracy across all training data fractions.
-Additionally, LLama3.2-1B maintains better stability as the training data is reduced, indicating its robustness in low-resource scenarios, whereas RoBERTa-base exhibits a steeper decline in performance with reduced data availability.
-This shows that LLMs are more stable, and instruction fine-tuning further enhances their performance and resilience to limited training data.
-5.2.2 Ablation 2: Fine-Tuning Techniques (Parameter-Efficient Methods)
-Fine-tuning LLMs such as LLama3.2-1B and multimodal models like LLama3.2-11B-vision-Instruct can be computationally expensive due to the need to update all parameters during training.
-To address this, parameter-efficient fine-tuning (PEFT) [84] methods like Low-Rank Adaptation (LoRA) and Prefix-Tuning aim to reduce computational overhead by updating only a small subset of model parameters.
-Full fine-tuning (FT) updates all model parameters during training, requiring high computational and memory resources but generally achieving the best performance.
-LoRA introduces low-rank trainable layers to the frozen pretrained weights, allowing significant parameter savings while maintaining performance.
-Prefix-Tuning appends trainable prefix embeddings to the input representations, leaving the main model frozen and thus reducing computational demands.
-For this experiment, we used LLama3.2-1B-Instruct for text-only tasks and LLama3.2-11B-vision-Instruct for multimodal tasks (text + image).
-The models were trained on 80% of the dataset, with 20% held out for evaluation.
-The evaluation metrics used are pecision, recall, F1 score, and accuracy.
-The goal was to evaluate if these parameter-efficient methods could achieve competitive performance compared to full fine-tuning.
-Table 7: Performance Comparison of Fine-Tuning Techniques for LLaMA 3.2 Models.
-Model Method F1 Score Accuracy Drop (%)
-LLama3.2-1B (Text) Full FT 83.1% 82.5% -
-LoRA 82.3% 81.8% 0.8%
-Prefix-Tuning 80.7% 80.5% 2.4%
-LLama3.2-11B (Text + Vision) Full FT 81.7% 83.0% -
-LoRA 81.0% 82.0% 0.7%
-Prefix-Tuning 80.2% 81.2% 1.5%
-Table 7 shows that full fine-tuning delivers the best performance across both text-only and multimodal tasks but comes at the cost of higher memory and computational requirements.
-LoRA achieves competitive performance with only a 0.7-0.8% drop in F1 score and accuracy, making it a highly practical alternative for resource-constrained environments.
-Prefix-tuning shows slightly more degradation, with a 1.5-2.4% drop in F1 score and accuracy, but remains viable when computational resources are extremely limited.
-These results demonstrate that PEFT methods can achieve near-parity with full fine-tuning while significantly reducing the computational overhead, making them suitable for deployment on edge devices or in scenarios with limited training infrastructure.
-5.3 Qualitative Analysis and Bias Identification
-In this section, we analyze potential biases in model predictions by examining how these models reason about certain statements.
-We elicit the chain-of-thought from LLMs and VLMs to determine whether their reasoning (1) relies on stereotypes or over-generalizations, or (2) correctly identifies or dismisses biases where relevant.
-We use our fine-tuned Llama3.2 ($1$B and $11$B), Mistral-v0.3-7B, and Phi3-vision model for generating predictions.
-5.3.1 Text-based Qualitative Analysis and Bias Identification
-Table 8 summarizes typical examples of each model’s responses.
-For each input statement, the model predicts whether there is Likely or Unlikely bias, then provides a brief rationale.
-We focus on two main aspects:
-  1. 1.
-    Misclassified Bias: Cases where the model either overestimates or underestimates the presence of bias due to stereotype-based reasoning or a failure to recognize marginalized perspectives.
-  2. 2.
-    Correct Classification: Cases where the model appropriately recognizes or dismisses bias based on the content and context of the statement.
-Model Predic Example and Reasoning Bias Observed
-Llama3.2-1B Bias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
-Reasoning (model): “Nurses rarely contribute to surgical innovations; doctors usually lead.”
-Issue: Stereotype-based reasoning. The model assumes nurses cannot be innovators, reinforcing traditional roles and overlooking individual competence.
-Llama3.2-1B Unbias Input: “The committee led by women executives proposed significant policy reforms.”
-Reasoning (model): “Focus on reforms; no inherent bias.”
-Issue: Fails to acknowledge or explore the possibility of gender-based biases in leadership perceptions; too dismissive.
-Llama3.2-11B Unbias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
-Reasoning (model): “There is no explicit bias. Nurses can contribute to innovative practices, and the statement acknowledges that.”
-Correct Classification: No stereotype is implied; the model recognizes nurses as capable innovators.
-Mistral-7B-v0.3 Bias Input: “The startup founded by women faced challenges securing venture capital funding.”
-Reasoning (model): “Historical funding gaps for women-led startups indicate bias.”
-Issue & Strength: Identifies systemic funding biases (correct), but may overgeneralize by implying all such startups experience the same bias.
-Mistral-7B-v0.3 Unbias Input: “The rural development program focused on uplifting marginalized communities.”
-Reasoning (model): “Development programs are generally well-intentioned; bias is unlikely.”
-Issue: Overlooks potential systemic biases in implementation or coverage, trivializing complexities of marginalized communities.
-Mistral-7B-v0.3 Unbias Input: “The community initiative led by indigenous elders improved local education outcomes.”
-Reasoning (model): “Recognizes indigenous leadership positively without implying stereotypes. No bias is evident.”
-Correct Classification: The model captures a constructive scenario and does not assume negative stereotypes about indigenous communities.
-Table 8: Binary bias classification (Bias/Unbias) with examples of both misclassified and correctly classified statements.
-Our observations in Table 8 reveal that some models (e.g., smaller Llama3.2 variants) occasionally rely on stereotypes (e.g., downplaying a nurse’s capability).
-Models often overlook subtle socio-cultural biases, especially when they assume that “no explicit mention of bias implies no bias”.
-In some cases, models correctly identify known systemic biases (e.g., women-led startups facing funding challenges) or offer neutral interpretations when appropriate.
-We also observe instances where models accurately identify or dismiss bias.
-For example, Llama3.2-11B correctly recognizes that highlighting a nurse’s innovative contributions is not inherently biased, and Mistral-7B-v0.3 appropriately identifies systemic gender biases for women-led startups.
-These cases show promise that with careful fine-tuning and context-awareness prompts, models can be guided to produce balanced and accurate assessments regarding bias.
-Overall, these qualitative findings help us refine the models’ training objectives, such as incorporating more nuanced real-world scenarios into the fine-tuning data and instructing models to critically assess context.
-5.3.2 Multimodal Qualitative Analysis and Bias Identification
-Model Input and Analysis
-Llama3.2-vision-11B
-Caption: “A program empowering indigenous women to advocate for their rights.”
-Reasoning: The text highlights empowerment, but the image lacks visible advocacy efforts, suggesting underrepresentation.
-Bias Observed: The model assumes that the absence of explicit visual markers is equal to bias, which may be an overgeneralization.
-Llama3.2-vision-11B
-Caption: “A collaborative community event highlighting diverse voices in science.”
-Reasoning: The text emphasizes diversity, but the image lacks representation of marginalized groups, indicating bias.
-Bias Observed: The model overrelies on textual implications of diversity and fails to interpret the broader visual context.
-Phi-3 Vision
-Caption: “A vibrant cultural festival in Africa.”
-Reasoning: The image does not convey the vibrancy described in the text, suggesting bias.
-Bias Observed: The model overemphasizes textual cues without fully considering the broader cultural context depicted in the image.
-Phi-3 Vision
-Caption: “A bustling shared street in Toronto during COVID.”
-Reasoning: While the image aligns with the description, certain social cues (e.g., accessibility) are missing, implying bias.
-Bias Observed: The model fails to account for nuanced social elements, such as inclusivity markers and diverse pedestrian usage.
-Llama3.2-vision-11B
-Caption: “Two individuals arrive at a runway for a formal event.”
-Reasoning: The individuals are dressed formally and wearing sunglasses at an airport. Nothing implies a stereotype or underrepresentation.
-Bias Observed: No bias observed. The model’s reasoning is correct and neutral—there is no mismatch between the caption and the image.
-Phi-3 Vision
-Caption: “A speaker waving to an audience at a tech conference.”
-Reasoning: The person is wearing business attire and appears to be greeting viewers. The caption is consistent with the visual context.
-Bias Observed: No bias observed. The model recognizes a straightforward scenario—no overgeneralization regarding the speaker’s background or role.
-Table 9: Binary classification results for VLMs. Images are displayed below model names to save horizontal space while maintaining readability.
-Table 9 illustrates how LLaMA3.2-vision-11B and Phi-3 Vision interpret combined textual and visual cues to classify potential bias. We identify two main tendencies:
-When a caption emphasizes marginalized groups or empowerment (e.g., “A program empowering indigenous women…”), LLaMA3.2-vision-11B often infers bias if the image lacks explicit advocacy symbols (e.g., protest signs), overgeneralizing beyond what is visually provided.
-A similar pattern emerges when discussing “diverse voices in science”, the model interprets the absence of visibly marginalized groups as bias without considering broader context.
-Phi-3 Vision sometimes focuses on literal textual descriptors (e.g., “vibrant cultural festival”) and deems any discrepancy as bias (e.g., the festival does not appear visually colorful).
-It may also note missing social cues (e.g., accessibility features) to imply bias, highlighting the difficulty of inferring inclusivity from a single snapshot.
-This can result in under-detection of subtler, systemic biases if the immediate visual–text match seems sufficient.
-Despite these tendencies, we also observe correct classifications when the text clearly aligns with the image:
-When the caption “A vibrant community gathering with multiple cultural performances” truly matches an image showing various cultural booths, attire, and families interacting, LLaMA3.2-vision-11B classifies bias as Unbias, noting “no sign of exclusion.” For a caption “An academic panel featuring speakers from multiple disciplines” paired with a visibly mixed group onstage, Phi-3 Vision detects “no explicit underrepresentation”, avoiding any unwarranted bias inference.
-These correct examples suggest that, with clearer text–image alignment and unambiguous markers of inclusivity, both models can produce balanced, context-sensitive judgments.
-However, ambiguous or subtle indicators of cultural norms or social empowerment remain challenging, which indicates a need for more context-aware multimodal reasoning.
+Figure 5: Top 10 outlets
+4 Results and Discussion
+4.1 Experimental Setting
+Compute & software. We run our experiments on NVIDIA A40/A100 Tensor Core GPUs. The stack used Python 3.10, PyTorch 2.0, Hugging Face Transformers, and OpenCV.
+Baselines and model families. We employed a suite of baseliens to evaluate on our data, which are:
+(1) Text-only SLMs: BERT-base-uncased [43], DistilBERT [44], RoBERTa-base [45], GPT-2 (small) [46], and BART-base [47].
+(2) Decoder-only LLMs (zero-/few-shot and IFT): open-source LLaMA-3.2-Instruct [2], Mistral-7B-Instruct [48]; closed-source: GPT-4o [40], Gemini 1.5 Pro [41].
+(3) VLMs (text+image): Phi-3-Vision-128k-Instruct [49], LLaVA-1.6 [50], LLaMA-3.2-11B-Vision-Instruct [2], plus GPT-4o-mini [40] and Gemini 1.5 Pro [41].
+For text–image annotation we used LLaVA-NeXT [51], Phi-3-Vision-128k-Instruct [52], and MiniCPM-V-2_6 [53], resolving conflicts via a two-step hierarchical voting scheme (details in Appx. C).
+Table 3: Vision–Language (Text+Image) Performance.
+Metrics are reported in percentages: Precision (Prec.), Recall, F1, and Accuracy (Acc.).
+Closed-source models are marked with .
+Results are grouped into SLMs and VLMs.
+For large models (Phi, LLaVA, LLaMA), we report Instruct variants.
+Configurations include fine-tuning (FT), instruction fine-tuning (IFT), and few-shot (0-shot, 5-shot).
+All models use the same splits and both modalities.
+Bold means best values in each model family.
+Model Config Prec. (%) Recall (%) F1 Score (%) Acc. (%)
+Small VLMs
+[…]
+[…]
+[…]
+[…]
+[…]
+Large VLMs
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Protocols and metrics.
+We evaluated models under three regimes: zero-shot, five-shot, and full fine-tuning.
+The dataset was partitioned into training, validation, and test splits with an 80/10/10 ratio, and we additionally employed 5-fold cross-validation to assess robustness.
+For classification tasks, we report standard metrics including precision, recall, F1 score, and accuracy.
+To capture performance on open-ended outputs, we further evaluate reasoning quality using LLM-as-judge assessments, focusing on reasoning accuracy and faithfulness.
+We used GPT4o as the judge model.
+We compare multimodal models in two complementary setups.
+In the closed-ended setting, models generate only a categorical label (biased or not biased), and evaluation relies on accuracy-based metrics against ground-truth annotations.
+In contrast, the open-ended setting (oVQA) requires models to produce both a label and a concise rationale.
+In this case, rationales are assessed by an LLM judge, which evaluates their correctness and consistency with the given evidence.
+Exploratory Analysis. Figure 5 presents an overview of our dataset through exploratory data analysis. More details of the dataset can be found in Appendix C Figure A.2.
+Table 4: Reasoning accuracy and faithfulness (LLM-judged) for open-ended rationales. Closed-source models are marked with $\dagger$}. Bold means best value in each model family.
+Model Config Reasoning Acc. (%) Faithfulness (%)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+4.2 Performance Evaluation
+In this section, we present the performance of multimodals on our dataset and show main results.
+4.2.1 Traditional fine-tuning is still competitive for classification
+We compare different VLMs (both small vs.
+large) and across regimes (0-shot, 5-shot, IFT) in a multimodal setting and present results in Table 3.
+The results show clear trends across scales and regimes.
+Among small VLMs, RoBERTa+CLIP attains the best F1 (82.9) and Acc (83.6), followed by BERT+CLIP (77.2 F1), highlighting gains from strong language encoders with multimodal backbones.
+Large open-source models improve from 0-/5-shot to IFT (e.g., Phi-3-Vision from 68.1 to 77.4 F1; LLaVA-1.6 from 61.6 to 75.0 F1).
+Closed-source systems perform strongly in 5-shot (e.g., GPT-4o mini 78.8 F1, Gemini-1.5 Pro 77.6 F1).
+Overall, our result findings show: (i) CLIP boosts small-model performance, (ii) IFT markedly improves large open-source models, and (iii) closed-source instruct models remain competitive.
+4.2.2 Open-ended reasoning (oVQA): Strong classifiers also tend to be strong Reasoners
+In the open-ended setting (oVQA), we evaluate autoregresive VLM models output both a label and a short rationale.
+We then use a judge LLM (GPT4o) to evaluate the rationale, reporting reasoning accuracy (the rate at which the rationale correctly supports the ground-truth decision) and faithfulness (the rate at which the rationale is supported by the given input).
+We also measure the correlation between closed-ended accuracy and open-ended reasoning accuracy across all models and regimes.
+6060656570707575808085859090505055556060656570707575808085859090Closed Accuracy (%)Reasoning Accuracy (%)Open-source VLMsClosed-source VLMs
+Figure 6: Closed vs open-ended reasoning accuracy across VLMs.
+The dashed diagonal indicates perfect alignment ($y=x$).
+Closed-source models (red triangles) cluster higher than open-source models (blue circles).
+The results in Table 4 show that reasoning accuracy is consistently lower than closed-ended accuracy, highlighting the difficulty of generating correct and grounded rationales compared to predicting labels.
+Across open-source models, reasoning accuracy ranges from 52–68% in the 0- and 5-shot settings, improving to nearly 69% under instruction tuning.
+Faithfulness scores follow a similar upward trend, with most models exceeding 75% once tuned, indicating that rationales are increasingly supported by the input.
+Closed-source systems (GPT-4o mini, Gemini-1.5 Pro) achieve higher reasoning accuracy and faithfulness overall, with GPT-4o IFT reaching close to 79% reasoning accuracy and nearly 89% faithfulness.
+Despite the performance gap as seen in Table 4, the correlation analysis (Pearson $r=0.91$) in Figure 6 shows a strong linear relationship between closed-ended accuracy and reasoning accuracy.
+This suggests that models which are strong classifiers also tend to be strong reasoners, although reasoning lags consistently behind classification by 6–12 percentage points.
+Instruction tuning narrows this gap for both open- and closed-source models, but does not eliminate it entirely.
+Overall, these result findings show that that closed-source VLMs are stronger reasoners, yet all models face a consistent reasoning gap compared to their closed-ended performance.
+4.2.3 Ablation Studies
+We analyze three factors that affect performance on text-only and multimodal tasks: (i) training data size, (ii) fine-tuning technique, and (iii) multimodal fusion strategy.
+Ablation 1: Training Data Size
+We examined how training data size affects performance across multiple VLMs: Phi-3-Vision, LLaVA-1.6, and LLaMA-3.2-11B-Vision. Each was trained on progressively larger subsets of the dataset.
+$10^{3.4}$$10^{3.6}$$10^{3.8}$$10^{4}$$10^{4.2}$$10^{4.4}$$10^{4.6}$$10^{4.8}$5050606070708080# Training Samples (log scale)F1 Score (%)Phi-3-VisionLLaVA-1.6LLaMA-3.2-11B-Vision
+Figure 7: Scaling curves of multiple VLMs trained with subsets of increasing size. All models improve with more data, though absolute F1 scores and scaling rates differ.
+Figure 7 compares three representative VLMs under different training data sizes.
+The results show that all models show consistent improvements as data grows, but with diminishing returns after the half-data mark.
+Phi-3-Vision achieves the highest overall F1 (82.0% at full scale), followed by LLaMA-3.2-11B-Vision (80.4%) and LLaVA-1.6 (78.8%).
+Interestingly, LLaVA lags at small scales but narrows the gap as more data is used, suggesting that weaker initial alignment can be partially compensated by scale.
+Across models, the relative gain from 1/16 to full data is 9–11 points, indicating that larger training sets are particularly beneficial for multimodal reasoning despite architectural differences.
+Ablation 2: Fine-Tuning Techniques (Parameter-Efficient Methods)
+We compare full fine-tuning against parameter-efficient methods across three representative VLMs: Phi-3-Vision, LLaVA-1.6, and LLaMA-3.2-11B-Vision.
+These techniques are recommended in the research to adjust a small subset of parameters (e.g., low-rank adapters, quantization-aware LoRA [54], lightweight adapters, or prompt embeddings) to reduce compute and memory costs while retaining model performance.
+Table 5: Comparison of parameter-efficient fine-tuning methods across three VLMs.
+Methods include: Full Fine-Tuning (FT), Low-Rank Adaptation (LoRA) [54], Quantized LoRA (QLoRA) [55], Adapters [56], and Prompt-Tuning [57].
+Bold highlights the best score per model.
+Trainable parameter percentages and GPU memory usage are shown for reference.
+[…]
+[…]
+[…]
+LoRA 1.2 10.1 81.2 66.8 77.9 64.7 79.7 65.3
+[…]
+Adapters 5.0 12.8 81.4 67.1 78.1 65.0 79.9 65.7
+[…]
+Table 5 shows that parameter-efficient methods achieve consistent trends across Phi-3-Vision, LLaVA-1.6, and LLaMA-3.2-11B-Vision.
+In all cases, LoRA and Adapters recover over 97–99% of the full fine-tuning performance while using less than 5% of trainable parameters.
+QLoRA offers nearly identical accuracy to LoRA but with lower memory requirements, confirming its utility for scaling.
+Prompt-tuning is the lightest-weight option but lags behind by 2–3 points in both F1 and reasoning accuracy.
+These results indicate that the efficiency–performance trade-off is robust across different multimodal backbones, with LoRA and QLoRA striking the most favorable balance.
+[…]
+Ablation 3: Performance gains from text-only to text+image modality
+We further quantify the impact of incorporating visual information by comparing text-only LLMs with their multimodal counterparts.
+For each model/regime, we pair the text-only variant with its closest text+image version and report the performance gap, defined as
+$\Delta=\mathrm{F1}_{\text{text+img}}-\mathrm{F1}_{\text{text}},$
+as shown in Figure 8.
+The results show that across ten paired models, the average gain is +8.22 F1 points (median +7.86; range +1.28+17.67).
+The largest improvements occur in smaller backbones and naïve fusion settings (e.g., DistilBERT+CLIP, $\Delta=+17.67$), whereas the gains diminish with stronger alignment and instruction tuning (e.g., LLaMA-3 family, IFT, $\Delta=+2.10$).
+Proprietary API results are excluded to avoid non-reproducible comparisons.
+Overall, these findings indicate that while text-only models are competitive, the addition of images consistently improves performance, particularly in challenging cases.
+Importantly, the performance gap narrows as multimodal alignment techniques mature, suggesting that visual grounding contributes most when models lack strong alignment mechanisms.
+Qualitative Analysis
+Figure 9: Sample examples from the our dataset showing the reasoning for the biased label.
+We present a qualitative analysis to illustrate how models reason about bias labels.
+A representative example is shown in Figure 9.
+We also conducted error analysis, and Appendix C Table A.2 demonstrates that models occasionally overgeneralize by flagging bias when captions reference marginalized groups or diversity but the accompanying image lacks explicit visual markers.
+Conversely, some models rely heavily on literal textual cues, treating minor mismatches as evidence of bias.
+While models generally perform well when textual and visual information are clearly aligned, ambiguous or nuanced cases continue to expose limitations in multimodal reasoning, underscoring the need for richer context modeling in bias detection.
 5 Discussion
 5.1 Practical and Social Impact
-Our findings carry both practical and social significance.
-By applying LLMs and VLMs, we introduce a scalable framework for detecting biases within multimodal media.
+Our findings carry both practical and social significance, as shown in Figure 10.
+By applying LLMs and VLMs, we introduce a scalable framework for detecting biases within multimodal news media.
 This is particularly relevant for content moderation, journalism, and public policy, where identifying subtle or underrepresented viewpoints can improve accountability and foster greater transparency.
@@
 In the long run, its widespread use could boost trust in digital content, broaden marginalized perspectives, and support fairer decision-making processes.
+Figure 10: Practical and social impact of ViLBias: a scalable LLM/VLM framework supports content moderation, journalism, and policy (practical), while advancing equity, transparency, and trust (social).
 5.2 Limitations
-Despite promising outcomes, our work comes with certain limitations.
-First, although our dataset spans a diverse range of media sources, it may not fully capture global cultural or linguistic nuances, which could affect the generalizability of our results.
-Second, our framework prioritizes biases that are more readily detected via textual or visual cues, potentially missing more pervasive systemic or cultural biases requiring deeper contextual insight.
-Additionally, dependence on SoTA language and vision-language models brings its own challenges, mainly significant computational overhead and resource requirements, which could be prohibitive for smaller organizations or those operating in lower-resource settings.
-While the hybrid human–machine annotation approach improves overall reliability, it is not entirely free from subjective disagreements, especially for ambiguous content [85].
-Media bias is also dynamic, shifting over time.
-Our dataset covers the period from May 2023 to September 2024, meaning newly emerging trends or evolving patterns may fall outside its scope.
-Future research should aim to include a broader range of cultures, languages, and emerging media practices, as well as more granular bias categories, while exploring real-time adaptability to keep pace with rapid media developments.
+Despite the strengths of our framework, several limitations remain.
+First, while our BiasCorpus spans a wide range of reputable outlets, its coverage is still skewed toward English-language and Western media.
+This restricts cross-cultural generalizability [58]and underrepresents media practices in regions such as Africa, South America, and South Asia.
+Similarly, the dataset employs binary labels (biased vs.
+not biased), which simplifies the inherently multi-dimensional nature of bias (e.g., partisan, sensationalist, representational).
+Extending to richer bias taxonomies remains an open challenge.
+Second, although our hybrid annotation pipeline mitigates noise through majority voting and HITL oversight, annotation quality is constrained by the interpretive subjectivity of bias judgments.
+Agreement between automated labels and human reviewers (Cohen’s $\kappa=0.72$) reflects substantial alignment but leaves room for disagreement, particularly in subtle or ambiguous cases.
+Moreover, because LLM as annotators may encode latent biases of their own [59], the benchmark risks inheriting systematic skew, raising the question of “bias in the bias benchmark’.’
+Third, our evaluation relies partly on LLM-as-judge assessments to score rationales.
+While efficient, this introduces circularity, since another model evaluates the faithfulness of generated explanations [42].
+Relatedly, reasoning judged as “faithful” may still omit deeper socio-cultural dynamics of framing, highlighting the gap between computational metrics and real-world fairness.
+Fourth, model and system constraints persist.
+Training and fine-tuning large VLMs remains computationally expensive, raising barriers for smaller labs and carrying environmental implications in line with ongoing Green AI debates [60].
+Furthermore, our benchmark does not explicitly test robustness under adversarial manipulations (e.g., image perturbations, paraphrased captions), which are increasingly relevant for deployment.
+Likewise, generalization to other domains such as health or social media memes remains unverified.
+Broader societal implications require caution.
+Automated bias detection tools are not value-neutral: if used irresponsibly by humans, they could reinforce rather than mitigate inequality.
+For example, governments or organizations may adopt such systems to suppress dissenting voices, while journalists or educators may over-rely on algorithmic judgments at the expense of critical human interpretation.
+These risks underscore the need for human accountability, transparency, and governance in every stage of system deployment [61].
+In addition, our framework does not yet address fairness across demographic stereotypes (e.g., race, gender, religion), an important future direction for ensuring that both humans and machines act responsibly in mitigating bias.
+In sum, ViLBias offers a scalable and rigorous benchmark, but its current scope, annotation design, and evaluation protocols should be seen as an initial step.
+Addressing cultural diversity, adversarial robustness, sustainability, and human responsibility will be essential for making multimodal bias detection truly comprehensive and socially beneficial.
 6 Conclusion
-In this study, we introduced ViLBias, a comprehensive framework designed to detect bias in multimodal contexts by capturing both linguistic and visual cues.
-Our results demonstrate that LLMs and VLMs can effectively identify subtle biases in text and images when supported by an intelligently curated dataset and a robust human–machine hybrid annotation strategy.
-Together, these efforts advance our goal of improving transparency in media and increasing trust in digital content.
-Performance metrics such as F1 score and accuracy confirm the reliability of LLM-based annotations.
-Furthermore, our majority-voting mechanism mitigates stochastic variations, enhancing the stability of the classification process.
-We also highlight the benefits of instruction-based fine-tuning and lightweight techniques like LoRA, enabling strong results even in resource-limited environments.
-Taken together, these contributions shows the potential of ViLBiasin bias detection in multimodal media scenarios.
-Declarations
-Funding: This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
-Conflict of Interest/Competing Interests: The authors declare that there are no conflicts of interest or competing interests associated with this study.
-Ethics Approval and Consent to Participate: Not applicable, as the study did not involve human participants or require ethical approval.
-Consent for Publication: All authors have reviewed and approved the final version of the manuscript and consent to its publication.
-Data Availability: The data supporting the findings of this study are available from the corresponding author upon reasonable request.
-Code Availability: The code used to implement the methods and experiments in this study is available upon publication.
-Author Contributions: S.R.: Writing – original draft, Validation, Methodology, Investigation, Conceptualization, Supervision.
-C.S.: Methodology, Investigation, Evaluation, Data Analysis.
-E.H.: Visualization, Methodology, Evaluation, Data Analysis.
-F.O.: Writing, Visualization, Methodology, Data Analysis.
-M.P.: Writing – review & editing, Methodology, Evaluation.
-V.C.: Methodology, Editing.
-R.J.: Methodology, Evaluation.
-M.L.: Data Curation.
-A.Z.: Writing – review & editing, Investigation.
-V.R.K.: Writing – review & editing, Investigation.
+In this work, we introduced ViLBias, a multimodal benchmark and framework for detecting and reasoning about bias in news media.
+By curating over 40,000 text–image pairs with hybrid LLM-as-annotator and HITL validation, we provide one of the first large-scale resources that evaluates both closed-ended classification and open-ended reasoning.
+Our experiments across SLMs, LLMs, and VLMs reveal consistent gains from multimodal grounding, demonstrate that parameter-efficient methods recover near full fine-tuning performance, and highlight the persistent gap between classification accuracy and reasoning quality.
+Together, these findings offer a rigorous foundation for advancing multimodal bias detection and evaluation.
+At the same time, our study underscores the importance of recognizing limitations.
+The dataset, while diverse, does not fully capture cultural and linguistic variation, and rationales remain shaped by subjective interpretations of bias.
+Evaluation protocols that rely on LLM-as-judges raise questions of circularity, while computational requirements restrict accessibility for smaller labs.
+Most importantly, the broader societal risks of misuse are non-trivial: if deployed irresponsibly by governments, media organizations, or educators, such systems could reinforce rather than mitigate inequality.
+Looking forward, we see three key directions.
+First, expanding coverage to more cultures, languages, and nuanced bias categories will strengthen representativeness [19].
+Second, robustness to adversarial manipulations and temporal shifts should be incorporated to ensure adaptability.
+Third, governance frameworks are needed to ensure that bias detection tools remain transparent, accountable, and augment rather than replace human judgment.
+By advancing technical capabilities while acknowledging risks and responsibilities, ViLBias takes a step toward fairer, more trustworthy multimodal media analysis.
 Appendix A Data Collection
@@
     Bias Detection: Identify any systematic biases, such as stereotypes or skewed representations in text or vision-text pairs, that might emerge from the model’s outputs.
-Annotators were specifically instructed to consider the following bias categories:
-    • Linguistic Biases: Patterns of stereotyping or unequal representation based on language usage, tone, or terminology.
-    • Visual Biases: Skewed interpretations or stereotypes inferred from visual elements.
-    • Gender Biases: Disparities or stereotypical roles related to gender in text or images.
-    • Racial and Ethnic Biases: Representations that reflect prejudice or unequal treatment based on race or ethnicity.
-    • Cultural Biases: Misinterpretations or insensitivity toward cultural norms or symbols.
-    • Disability Biases: Exclusion or misrepresentation of individuals with disabilities.
-    • Age Biases: Stereotyping or underrepresentation based on age.
+Annotators were specifically instructed to consider the following bias categories, such as linguistic, visual, gender, racial, cutlural , disbility and age related biases.
   6. 6.
@@
 The feedback collected was used to refine model prompts and guide further improvements in the annotation pipeline.
-Prompt for Text-based annotation prompt = f””” Analyze the provided text for evidence of bias.
+Appendix C Prompts
+Prompt for Text-based Annotation prompt = f""" Analyze the provided text for evidence of bias.
 Identify the presence or absence of rhetorical techniques that contribute to bias in the news content.
@@
 Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation concise and structured, without additional explanations.
-”””
-Prompt for Text and image based annotation prompt = f””” Analyze the provided text and associated image for bias.
+"""
+Prompt for Text and Image-based Annotation prompt = f""" Analyze the provided text and associated image for bias.
 Evaluate how the text and image interact to reinforce or contradict the narrative and identify techniques used to promote a specific perspective.
@@
 Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices or Visual Cues: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Symbolism: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation precise and avoid unnecessary explanations.
-”””
+"""
+LLM-as-Judge Prompt (ViLBias metrics: Reasoning Accuracy & Faithfulness) You are an impartial judge.
+Score a model’s open-ended rationale for a news text+image item on two binary metrics: (A) Reasoning Accuracy — does the rationale, if taken as true, correctly support the ground-truth label?
+(B) Faithfulness — is the rationale grounded in the provided input (text and/or image) with no hallucinated facts.
+Inputs - QUESTION: “Does this content exhibit bias?
+Please explain your reasoning.” - TEXT: <<<{TEXT SNIPPET}>>> - IMAGE: (rendered alongside; if unavailable, write “none”) - GROUND_TRUTH_LABEL: {biased | not_biased} - MODEL_PRED_LABEL: {biased | not_biased} - MODEL_RATIONALE: <<<{FREE-TEXT RATIONALE}>>> Judging Rules 1) Reasoning Accuracy (0/1): Assign 1 if the rationale’s claims, taken at face value, logically entail or strongly support the GROUND_TRUTH_LABEL.
+Penalize if it argues for the opposite label, is purely label restatement with no reason, or uses irrelevant logic.
+2) Faithfulness (0/1): Assign 1 only if the rationale’s cited cues are verifiably present in TEXT and/or IMAGE.
+Disallow unstated assumptions, world knowledge, or fabricated details.
+When possible, quote minimal text spans and/or describe visual evidence (objects/attributes/relations); if no image, judge solely on text.
+3) Edge Cases: - Empty, generic, or circular rationales (“it is biased because it is biased”) $\rightarrow$ ReasoningAccuracy=0, Faithfulness=0.
+- Correct label but incorrect/unsupported reasons $\rightarrow$ ReasoningAccuracy=0 or Faithfulness=0 as appropriate.
+- If MODEL_PRED_LABEL differs from GROUND_TRUTH_LABEL, you still judge the rationale against the ground truth.
+Output (strict JSON)
+  "reasoning_accuracy": 0 or 1,
+  "faithfulness": 0 or 1,
+  "justification": "2-3 sentences: why each score was assigned"
+Grading Hints - Typical textual bias cues: loaded wording, one-sided framing, selective emphasis/omission.
+- Typical visual bias cues: image choice, cropping/staging, emotive imagery, text–image mismatch.
+- Keep judgments short, precise, and grounded only in the provided inputs.
+Figure A.1: LLM-as-Judge prompt for evaluation
+Table A.1: Key settings for training, generation, and evaluation.
+Generation Temperature 0.2–0.4; max tokens 1024
+Batch / LR BERT: 32, 15 epochs, $2\!\times\!10^{-5}$, wd 0.01, ES patience 5
+FT (VLM/LLM): batch 1, 3 epochs, $2\!\times\!10^{-4}$, grad acc 1, max grad norm 0.3
+Scheduler Constant LR with 0.03 warmup ratio
+Precision Mixed precision (bf16=True)
+Loss CrossEntropy; Focal ($\alpha{=}1,\ \gamma{=}2$) where noted
+Eval Accuracy, Precision, Recall, F1; confusion matrices (batch 32)
+Quant. 4-bit (nf4) with double quantization
+Reproduc. CUDA devices (CPU fallback), seed 0; TOKENIZERS_PARALLELISM=false
+Table A.2: Binary classification sample results for various VLMs.
+Model Input and Analysis
+Llama3.2-vision-11B
+Caption: “A program empowering indigenous women to advocate for their rights.”
+Reasoning: The text highlights empowerment, but the image lacks visible advocacy efforts, suggesting underrepresentation.
+Bias Observed: The model assumes that the absence of explicit visual markers is equal to bias, which may be an overgeneralization.
+Llama3.2-vision-11B
+Caption: “A collaborative community event highlighting diverse voices in science.”
+Reasoning: The text emphasizes diversity, but the image lacks representation of marginalized groups, indicating bias.
+Bias Observed: The model overrelies on textual implications of diversity and fails to interpret the broader visual context.
+Phi-3 Vision
+Caption: “A vibrant cultural festival in Africa.”
+Reasoning: The image does not convey the vibrancy described in the text, suggesting bias.
+Bias Observed: The model overemphasizes textual cues without fully considering the broader cultural context depicted in the image.
+Phi-3 Vision
+Caption: “A bustling shared street in Toronto during COVID.”
+Reasoning: While the image aligns with the description, certain social cues (e.g., accessibility) are missing, implying bias.
+Bias Observed: The model fails to account for nuanced social elements, such as inclusivity markers and diverse pedestrian usage.
+Llama3.2-vision-11B
+Caption: “Two individuals arrive at a runway for a formal event.”
+Reasoning: The individuals are dressed formally and wearing sunglasses at an airport. Nothing implies a stereotype or underrepresentation.
+Bias Observed: No bias observed. The model’s reasoning is correct and neutral—there is no mismatch between the caption and the image.
+Phi-3 Vision
+Caption: “A speaker waving to an audience at a tech conference.”
+Reasoning: The person is wearing business attire and appears to be greeting viewers. The caption is consistent with the visual context.
+Bias Observed: No bias observed. The model recognizes a straightforward scenario—no overgeneralization regarding the speaker’s background or role.
+(a) Categories covered by some news outlets. Color represents bias. Size of marker represents the number of articles which have the same newspaper, category and bias.
+(b) News outlet and the ratio of biased articles that are released. Size of marker represents sample size of articles for that news outlet.
+(c) Outlet-wise distribution of bias for each category. Size of marker represents sample size. Vertical position represents the amount of bias of a newspaper in that category.
+Figure A.2: Distribution charts of bias counts and ratios by outlets and categories.